### v1

# Jacobi’s solution for geodesics on a triaxial ellipsoid

Abstract
On Boxing Day, 1838, Jacobi found a solution to the problem of geodesics on a triaxial ellipsoid, with the course of the geodesic and the distance along it given in terms of one-dimensional integrals. Here, a numerical implementation of this solution is described. This entails accurately evaluating the integrals and solving the resulting coupled system of equations. The inverse problem, finding the shortest path between two points on the ellipsoid, can then be solved using a similar method as for biaxial ellipsoids.
geometrical geodesy, geodesics, triaxial ellipsoids, numerical methods
1 Introduction
The geodesic is the shortest path between two points on a surface, and it plays a crucial role in geodesy, where the earth is typically modeled as an ellipsoid of revolution, a biaxial ellipsoid. The main geodesic problems are (1) given a starting point and a direction, find the point a certain distance away, the “direct” geodesic problem, and (2) to find the length and direction of the geodesic connecting two points, the “inverse” geodesic problem. The path of a geodesic is also given by the motion of a point mass sliding on the surface of the ellipsoid without friction and in the absence of external forces. This allows geodesics to be extended indefinitely.
Recently, there has been interest in using a refined model of the earth as a triaxial ellipsoid (Panou et al., 2020). Consequently, there is interest in obtaining reliable solutions to the geodesic problem in this case.
The solution of the direct geodesic problem in the biaxial case is relatively straightforward. The angular momentum about the axis of symmetry (the Clairaut constant) is conserved, allowing the path to be found in terms of elliptic integrals.
The triaxial ellipsoid, on the other hand, possesses no obvious symmetry. It therefore came as a surprise when Jacobi (1839) found that the geodesic problem could be reduced to quadrature in this case too, with the solution given in terms of one-dimensional integrals. (We know the date, given in the abstract, for this discovery, because of a letter he wrote on December 28, 1838, to F. W. Bessel, his neighbor in Königsberg.) Jacobi (1843, §28) expanded on his method in his Lectures on Dynamics, and the result was generalized by Liouville (1846, §§20–21) to apply to so-called Liouville surfaces. The qualitative properties of the solution can be found in several textbooks (Darboux, 1894; Hilbert and Cohn-Vossen, 1952; Klingenberg, 1982; Arnol’d, 1989; Berger, 2010). However, unlike the case of the biaxial ellipsoid, where Bessel (1825) provided a prescription for computing geodesics, little effort was given to implementing Jacobi’s solution.
The goal of this paper is to address this deficiency, specifically to enable the direct and inverse problems to be solved with high accuracy and reasonable efficiency. This entails approximating the integrands as Fourier series, which allows the indefinite integrals to be easily evaluated, and finding an efficient way to solve the resulting coupled system of equations. With the solution to the direct problem in hand, we turn to solving the inverse problem, following the same basic recipe used in the biaxial case (Karney, 2013, 2024a).
Panou (2013); Panou and Korakitis (2019) explore an alternative approach to solving the direct geodesic problem, namely by numerically integrating the corresponding ordinary differential equations for the geodesics, as discussed in Appendix E. This can provide an accurate solution, although the properties of the true solution are only approximately maintained. They do not provide a complete solution to the inverse problem.
2 Ellipsoidal coordinates
[…]
[scale=0.75]figures/graticule
Figure 1: The ellipsoidal grid showing lines of constant $\beta$ and $\omega$. The grid spacing is $15^{\circ}$. The heavy lines show the minor ($X=0$ or $\cos\omega=0$), median ($Y=0$ or $\cos\beta\sin\omega=0$), and major ($Z=0$ or $\sin\beta=0$) principal ellipses of the ellipsoid. The parameters of the ellipsoid are $a=1.01$, $b=1$, and $c=0.8$, and it is viewed in an orthographic projection looking at the point with geodetic coordinates $\phi=40^{\circ}$, $\lambda=30^{\circ}$.
Jacobi’s insight was to express the equations of the geodesic in terms of ellipsoidal coordinates; this allows the equations to be reduced to one-dimensional integrals through the separation of variables.
Consider the ellipsoid defined by
$S(\mathbf{R})=\frac{X^{2}}{a^{2}}+\frac{Y^{2}}{b^{2}}+\frac{Z^{2}}{c^{2}}-1=0,$ (1)
where $\mathbf{R}=[X,Y,Z]^{T}$ is a three-dimensional point, and $a$, $b$, and $c$ are the major, median, and minor semiaxes, satisfying $a\geq b\geq c>0$. (The superscript $T$ means “transpose,” converting a row vector into a column vector.) We characterize the shape of the ellipsoid by the parameters,
$e=\frac{\sqrt{a^{2}-c^{2}}}{b},\quad k=\frac{\sqrt{b^{2}-c^{2}}}{\sqrt{a^{2}-c^{2}}},\quad k^{\prime}=\frac{\sqrt{a^{2}-b^{2}}}{\sqrt{a^{2}-c^{2}}}.$ (2)
Here $e$ measures how much the ellipsoid departs from a sphere, while $k$ and $k^{\prime}$ describe how close the ellipsoid is to being oblate ($k=1$) or prolate ($k=0$); note that $k^{2}+k^{\prime 2}=1$. The semiaxes are related to these parameters by
$a,b,c]=b\bigl[\sqrt{1+e^{2}k^{\prime 2}},1,\sqrt{1-e^{2}k^{2}}\bigr].$ (3)
The case $c=0$, where the ellipsoid becomes an elliptical disc, is briefly discussed in Appendix D.
A point on the ellipsoid can be written in terms of ellipsoidal coordinates, the latitude, $\beta$, and the longitude, $\omega$, as
$\mathbf{R}=\begin{bmatrix}a\cos\omega\sqrt{k^{2}\cos^{2}\beta+k^{\prime 2}}\\ b\cos\beta\sin\omega\\ c\sin\beta\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\end{bmatrix}$. (4)
Lines of constant $\beta$ and $\omega$ define a grid, illustrated in Fig. 1. The grid lines of the ellipsoidal coordinates are “lines of curvature” on the ellipsoid, i.e., they are parallel to the directions of principal curvature. The coordinates are singular at the umbilics, $\cos\beta=\sin\omega=0$, where the principal curvatures are equal (locally, the ellipsoid is spherical). The grid lines are also intersections of the ellipsoid with confocal systems of hyperboloids of one and two sheets. Finally, the lines of curvature are geodesic ellipses and hyperbolas, where the foci are neighboring umbilics.
In the limit $k\rightarrow 1$ (resp. $k\rightarrow 0$), the umbilics converge on the $Z$ (resp. $X$) axis and an oblate (resp. prolate) ellipsoid is obtained with $\beta$ (resp. $\omega$) becoming the standard parametric latitude and $\omega$ (resp. $\beta$) becoming the standard longitude. The sphere is a non-uniform limit, with the position of the umbilics depending on $k$.
Define three vectors giving the “East,” “North,” and “Up” directions:
$\mathbf{E}$ $=\partial\mathbf{R}/\partial\omega=\begin{bmatrix}-a\sin\omega\sqrt{k^{2}\cos^{2}\beta+k^{\prime 2}}\\[2.15277pt] b\cos\beta\cos\omega\\[2.15277pt] \displaystyle c\frac{k^{\prime 2}\sin\beta\cos\omega\sin\omega}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}}\end{bmatrix}$, (5a)
$\mathbf{N}$ $=\partial\mathbf{R}/\partial\beta=\begin{bmatrix}\displaystyle-a\frac{k^{2}\cos\beta\sin\beta\cos\omega}{\sqrt{k^{2}\cos^{2}\beta+k^{\prime 2}}}\\[9.47217pt] -b\sin\beta\sin\omega\\[2.15277pt] c\cos\beta\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\end{bmatrix}$, (5b)
$\mathbf{U}$ $={\textstyle\frac{1}{2}}\nabla S(\mathbf{R})=\biggl[\frac{X}{a^{2}},\frac{Y}{b^{2}},\frac{Z}{c^{2}}\biggr]^{T}.$ (5c)
It is easy to verify that $\mathbf{N}\cdot\mathbf{E}=0$, so that $\mathbf{E},\mathbf{N},\mathbf{U}]$ are mutually orthogonal. As a consequence, the element of distance $\mathrm{d}s$ for the ellipsoidal coordinate system is given by
$\frac{\mathrm{d}s^{2}}{b^{2}}$ $=\frac{\left|\mathbf{N}\right|^{2}\,\mathrm{d}\beta^{2}+\left|\mathbf{E}\right|^{2}\,\mathrm{d}\omega^{2}}{b^{2}}$
$=(k^{2}\cos^{2}\beta+k^{\prime 2}\sin^{2}\omega)$
$\quad{}\times\biggl(\frac{1-e^{2}k^{2}\cos^{2}\beta}{k^{\prime 2}+k^{2}\cos^{2}\beta}\,\mathrm{d}\beta^{2}+\frac{1+e^{2}k^{\prime 2}\sin^{2}\omega}{k^{2}+k^{\prime 2}\sin^{2}\omega}\,\mathrm{d}\omega^{2}\biggr).$ (6)
Furthermore, the direction of a geodesic is
$\mathbf{V}=\sin\alpha\hat{\mathbf{E}}+\cos\alpha\hat{\mathbf{N}}$, (7)
where $\alpha$ is the azimuth of the geodesic measured clockwise from a line of constant $\omega$. At the pole of an oblate ellipsoid, we take the limit $\cos\beta\rightarrow 0+$, to give
$\hat{\mathbf{E}}$ $=[-\sin\omega,\cos\omega,0]^{T},$ (8a)
$\hat{\mathbf{N}}$ $=\sin\beta[-\cos\omega,-\sin\omega,0]^{T}.$ (8b)
Similarly, at the pole of a prolate ellipsoid, we take the limit $\sin\omega\rightarrow 0+$, to give
$\hat{\mathbf{E}}$ $=\cos\omega[0,\cos\beta,\sin\beta]^{T},$ (9a)
$\hat{\mathbf{N}}$ $=[0,-\sin\beta,\cos\beta]^{T}.$ (9b)
At an umbilic on a general ellipsoid, $\cos\beta\rightarrow 0$ and $\sin\omega\rightarrow 0$, we have $\left|\mathbf{E}\right|=\left|\mathbf{N}\right|=0$ so that $\hat{\mathbf{E}}$ and $\hat{\mathbf{N}}$ become ill-defined. In this case, we use the conventional geodetic definitions of $\hat{\mathbf{E}}$ and $\hat{\mathbf{N}}$,
$\hat{\mathbf{U}}$ $=[ck^{\prime}\cos\omega,0,ak\sin\beta]^{T}/b,$ (10a)
$\hat{\mathbf{E}}$ $=[0,\cos\omega,0]^{T},$ (10b)
$\hat{\mathbf{N}}$ $=\hat{\mathbf{U}}\times\hat{\mathbf{E}}$. (10c)
For geodesics that intersect an umbilic, we have
$\tan\alpha=\pm\frac{k^{\prime}}{k}\frac{\sin\omega}{\cos\beta};$ (11)
this follows from setting $\gamma=0$ in Eq. (14), given below. Expanding $\mathbf{R}$ about an umbilic to second order in $\cos\beta$ and $\sin\omega$, we find
[…]
where we have chosen the sign in Eq. (11) as $\pm 1=-\sin\beta\cos\omega$ to yield the normal convention that $\alpha$ measures angles clockwise.
The torus $(\omega,\beta)\in[-\pi,\pi]\times[-\pi,\pi]$ covers the ellipsoid twice. To facilitate passing to the limit of an oblate ellipsoid, we may regard $\pi,\pi]\times[-{\textstyle\frac{1}{2}}\pi,{\textstyle\frac{1}{2}}\pi]$ as the principal sheet and insert branch cuts at $\beta=\pm{\textstyle\frac{1}{2}}\pi$. The rule for switching sheets is
$\omega\rightarrow-\omega,\quad\beta\rightarrow\pi-\beta,\quad\alpha\rightarrow\pi+\alpha.$ (13)
Other coordinate systems are frequently used for an ellipsoid: geodetic, parametric, and geocentric. Conversions between the various coordinate systems are considered in Appendix A.
3 Qualitative behavior
[…]
[scale=0.75]figures/geodesicc
[…]
Figure 2: Samples of geodesics on an ellipsoid with the same parameters and viewpoint as Fig. 1. The parameters of the geodesics are given in Table 1. These figures are adapted from figures that the author contributed to Wikipedia (Wikipedia contributors, 2013).
Table 1: Parameters for the sample geodesics shown in Fig. 2. The starting points are given by $\beta_{1}$, $\omega_{1}$, and $\alpha_{1}$. The corresponding value of $\gamma$ is given by Eq. (14). The initial conditions are such that the starting points lie on the median principal ellipse $Y=0$ and the initial direction is $\mathbf{V}_{1}=[0,1,0]^{T}$. The geodesics are followed a distance $\pm s_{12}$ in each direction. The ratio $r=p/q$ indicates that, over the full distance $2s_{12}$, the geodesic executes $p$ complete oscillations/circuits in $\beta$ and $q$ circuits/oscillations in $\omega$.
$\beta_{1}\,(^{\circ})$ $\omega_{1}\,(^{\circ})$ $\alpha_{1}\,(^{\circ})$ $\gamma$ $s_{12}$ $r$
(a) $42.70330$ 0 $90$ $0.51148$ $162.80637$ $61/52$
(b) $87.52250$ 0 $90$ $0.00177$ $247.24408$ $87/85$
(c) $90$ 0 $135$ 0 $142.63587$ $50/50$
(d) $90$ $10.15216$ $180$ 0.00164-0.00164 $252.96477$ $89/87$
(e) $90$ $39.25531$ $180$ 0.02117-0.02117 $156.05191$ $55/53$
Let us illustrate the qualitative properties of geodesics; these are readily found from the form of Jacobi’s solution and are described in the textbooks listed in the introduction. On a given geodesic,
$\gamma=k^{2}\cos^{2}\beta\sin^{2}\alpha-k^{\prime 2}\sin^{2}\omega\cos^{2}\alpha$ (14)
is a constant. This is a generalization of the familiar Clairaut constant, which characterizes geodesics on a biaxial ellipsoid. Figure 2 shows samples of geodesics on an ellipsoid with $a=1.01$, $b=1$, and $c=0.8$ (the same parameters as Fig. 1); the values of $\gamma$ for these geodesics are given in Table 1. Ignoring for now Fig. 2(c), we see that, depending on whether $\gamma$ is positive or negative, either $\omega$ or $\beta$ is a “rotating” coordinate (increasing or decreasing without limit), and, correspondingly, $\beta$ or $\omega$ is a “librating” coordinate (oscillating about a fixed value). We label these two cases “circumpolar”, $\gamma>0$, Figs. 2(a, b), and “transpolar”, $\gamma<0$, Figs. 2(d, e).
The circumpolar geodesics are similar to the geodesics on an oblate ellipsoid and a limiting case of such geodesics is the major principal ellipse defined by $Z=0$. Likewise, the transpolar geodesics mimic the geodesics on a prolate ellipsoid, and a limiting case of such geodesics is the minor principal ellipse defined by $X=0$. The transition between these two classes of geodesics is shown in Fig. 2(c), where $\gamma=0$. In this case, the geodesic—an umbilical geodesic—repeatedly crosses two opposite umbilics; following the geodesic in either direction, it eventually lies on the median principal ellipse $Y=0$.
For a biaxial ellipsoid, the equator and all the meridians are simple (not self-intersecting) closed geodesics. On the other hand, for a triaxial ellipsoid (provided it is not too eccentric), there are only three simple closed geodesics, namely the 3 principal ellipses. The major and minor ellipses are stable; if they are perturbed, the resulting geodesic oscillates about the original ellipse. However, the median ellipse is unstable; if the geodesic is perturbed, it swings away from the $Y=0$ plane before returning to the original ellipse, but now traveling along it in the opposite direction. The stability of closed geodesics is treated in Appendix F.
For almost all $\gamma\neq 0$, a geodesic covers the area bounded by the limiting lines of curvature. The examples of geodesics in Figs. 2(a, b, d, e) are exceptional in that they are closed. For the corresponding values of $\gamma\neq 0$ listed in Table 1, the geodesics are closed in the same way regardless of the initial conditions. For a particular such $\gamma$, the union of the closed geodesics is area filling; this is an example of Poncelet’s porism. Umbilical geodesics, exemplified by Fig. 2(c), are not area filling, but here again, the union of all such geodesics is, covering the entire ellipsoid.
4 Jacobi’s solution
[…]
[scale=0.75]figures/cayley
Figure 3: The graphical method for plotting umbilical geodesics given by Cayley (1872). The semiaxes of the ellipsoid are $a=\sqrt{1000}$, $b=\sqrt{500}$, $c=\sqrt{250}$, and it is viewed here in an orthogonal projection along the $Y$ axis. The dotted lines are lines of $\beta=\beta^{(f)}_{n}$ (going bottom to top) or $\omega=\omega^{(f)}_{n}$ (going left to right), as defined in the text, with the separation constant given by $\Delta^{(f)}=1/\sqrt{160}$ (matching Cayley’s choice). In this figure, the lines of constant $\beta$ and $\omega$ are labeled with the corresponding values of $n$ (these values are given in Table 2). The geodesics, shown as heavy lines, connect the vertices of the resulting mesh, and all converge on the umbilic labeled $U$, or on the neighboring one (not shown in the figure).
[…]
[scale=0.75]figures/cayleydist
Figure 4: Marking the distance along geodesics. The heavy lines are those geodesics in Fig. 3 that converge on the umbilic $U$. As in that figure, the dotted lines are lines of constant $\beta$ or $\omega$; however, in this case, the values are given by equal increments of $\Delta^{(g)}=1/10$ in the distance functions (these values are given in Table 2). The dashed lines connect the vertices of the mesh, and these mark off distance intervals of $b/10$ along the geodesics.
Here we summarize the solution of the geodesic problem following Darboux (1894, §§583–585); a comparable treatment is given by Klingenberg (1982, §§3.5.4–3.5.6). The expression for $\mathrm{d}s^{2}$ Eq. (6) fulfills the condition of a “Liouville surface,” with metric given by Darboux’s Eq. (23),
$\mathrm{d}s^{2}=\bigl(U-V\bigr)\bigl(U_{1}^{2}\,\mathrm{d}u^{2}+V_{1}^{2}\,\mathrm{d}v^{2}\bigr),$ (15)
where $U$ and $U_{1}$ are functions of $u$ and $V$ and $V_{1}$ are functions of $v$. Identifying
$(u,v)$ $=(\beta,\omega),$ (16a)
$(U,V)$ $=(k^{2}\cos^{2}\beta,-k^{\prime 2}\sin^{2}\omega),$ (16b)
$(U_{1}^{2},V_{1}^{2})$ $=\biggl(\frac{1-e^{2}k^{2}\cos^{2}\beta}{k^{\prime 2}+k^{2}\cos^{2}\beta},\frac{1+e^{2}k^{\prime 2}\sin^{2}\omega}{k^{2}+k^{\prime 2}\sin^{2}\omega}\biggr),$ (16c)
the course of the geodesic is given by Darboux’s Eq. (28),
$\delta$ $=\int\frac{\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
$\qquad{}\mp\int\frac{\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega,$ (17a)
and the distance $s$ along the geodesic is given by Darboux’s Eq. ($33^{\prime}$),
$\frac{s+s_{1}}{b}$ $=\int\frac{k\cos^{2}\beta\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
$\quad{}\pm\int\frac{k^{\prime}\sin^{2}\omega\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega.$ (17b)
Here $\delta$ and $s_{1}$ are constants given by the initial conditions. Except at umbilics, the direction of the line is determined by the constant $\gamma$, defined in Eq. (14) and given by Darboux’s Eq. (30). At umbilics, $\gamma$ vanishes, and the direction is given by $\delta$.
The integrals in Eqs. (17) are related to one another. It is therefore convenient to define
$f(\phi;\kappa,\epsilon,\mu)$ $=\int_{0}\frac{\sqrt{1-\epsilon\kappa\cos^{2}\phi}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\phi}\sqrt{\kappa\cos^{2}\phi+\mu}}\,\mathrm{d}\phi,$ (18a)
$g(\phi;\kappa,\epsilon,\mu)$ $=\int_{0}\frac{\kappa\cos^{2}\phi\sqrt{1-\epsilon\kappa\cos^{2}\phi}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\phi}\sqrt{\kappa\cos^{2}\phi+\mu}}\,\mathrm{d}\phi,$ (18b)
where $\kappa\in[0,1]$, $\kappa^{\prime}=1-\kappa$, $\mu\in[-\kappa;\kappa^{\prime}]$, $\epsilon\in(-\infty,1/\kappa)$. Equations (17) can be written as
$\delta$ $=f(\beta;k^{2},e^{2},-\gamma)$
$\qquad{}\mp f(\omega-{\textstyle\frac{1}{2}}\pi;k^{\prime 2},-e^{2},\gamma),$ (19a)
$(s+s_{1})/b$ $=g(\beta;k^{2},e^{2},-\gamma)$
$\qquad{}\pm g(\omega-{\textstyle\frac{1}{2}}\pi;k^{\prime 2},-e^{2},\gamma).$ (19b)
At present, we leave the signs of the square roots in the integrals unspecified. However, the presence of $\mp$ and $\pm$ in these equations indicates that while progressing along a geodesic, the two terms in Eq. (19a) cancel while those in Eq. (19b) combine. In the following, we drop the parametric arguments for the $f$ and $g$ functions; $\kappa,\epsilon,\mu]=[k^{2},e^{2},-\gamma]$ are implied for functions of $\beta$, and $k^{\prime 2},-e^{2},\gamma]$ for functions of $\omega$.
Table 2: Values of $\beta^{(f)}_{n}$, $\omega^{(f)}_{n}$, $\beta^{(g)}_{n}$, and $\omega^{(g)}_{n}$ used for the coordinate meshes in Figs. 3 and 4. For the coordinates in the 2nd and 3rd columns, the values of the $f$ functions are multiples of $\Delta^{(f)}=1/\sqrt{160}$, and the coordinates are used in Fig. 3. For those in the 4th and 5th columns, the values of the $g$ functions are multiples of $\Delta^{(g)}=1/10$, and the coordinates are used in Fig. 4.
[…]
0 $\hphantom{0}0$ $\hphantom{0}90$ $\hphantom{0}0$ $\hphantom{0}90$
11 $\hphantom{0}7.789$ $\hphantom{0}95.538$ $13.993$ $\hphantom{0}94.967$
22 $15.265$ $101.015$ $27.852$ $\hphantom{0}99.966$
33 $22.205$ $106.377$ $41.915$ $105.029$
44 $28.511$ $111.571$ $57.515$ $110.195$
55 $34.175$ $116.557$ $84.901$ $115.507$
66 $39.237$ $121.301$ $121.023$
77 $43.758$ $125.782$ $126.818$
88 $47.803$ $129.985$ $133.004$
99 $133.906$ $139.762$
$10$ $54.696$ $137.547$ $147.434$
$11$ $140.916$ $156.870$
$12$ $60.309$ $144.024$ $173.205$
$14$ $149.518$
$16$ $68.769$ $154.156$
2020 $161.362$
$30$ $171.645$
The structure of Eqs. (19a) allows geodesics to be traced by a simple construction given by Cayley (1872), who considered umbilical geodesics, $\gamma=0$, on an ellipsoid with $a:b:c=\sqrt{2}:1:1/\sqrt{2}$. Find the values of $\beta$ (resp. $\omega$), such that $f(\beta^{(f)}_{n})=n\Delta^{(f)}$ (resp. $f({\omega^{(f)}_{n}-{\textstyle\frac{1}{2}}\pi})=n\Delta^{(f)}$). Now draw the grid lines $\beta=\beta^{(f)}_{n}$ and $\omega=\omega^{(f)}_{n}$, forming a mesh on the ellipsoid. Two families of geodesics can be traced through the mesh by connecting opposite corners of each cell. This follows from Eq. (19a), and the method only “works” because of the separation of variables in the solution. The result is shown in Fig. 3, where we reproduce the case examined by Cayley (1872, Plate II, following p. 130). (Besides some understandable errors arising from the low-order methods he used to evaluate and invert the integrals, Cayley made some mistakes connecting the vertices of the mesh.)
The same construction can be used to mark off distances along the geodesics using Eq. (19b). In this case, we construct a mesh defined by $g(\beta^{(g)}_{n})=n\Delta^{(g)}$ and $g({\omega^{(g)}_{n}-{\textstyle\frac{1}{2}}\pi})=n\Delta^{(g)}$ as shown in Fig. 4. We see that the solution of the direct geodesic problem essentially reduces to “tabulating” four one-dimensional integrals.
5 Properties of general geodesics
The course of a geodesic is determined by Eq. (19a) and we are immediately confronted with the problem that, for $\gamma\neq 0$, one of the integrands in Eq. (17) is singular and that, if $\gamma=0$, the integrals themselves are singular. In this section, we address the circum- and transpolar cases where $\gamma\neq 0$.
We deal with these cases together by denoting the rotating and librating coordinates as $\theta$ and $\phi$, respectively. We include the offset of ${\textstyle\frac{1}{2}}\pi$ in relations involving $\omega$. Let us also define $\tau$ as the azimuth measured from a line of constant $\theta$. Because $\phi$ and $\tau$ are librating coordinates, it’s useful to introduce constants $S_{\phi}=\operatorname{sign}(\cos\phi)$ and $S_{\tau}=\operatorname{sign}(\sin\tau)$ to specify the values about which $\phi$ and $\tau$ oscillate.
For the rotating coordinate $\theta$, we fold in the direction of the geodesic so that $\theta$ increases in the forward direction. Thus, for circumpolar geodesics, we have
$\phi=\beta,\quad\tau=\alpha,\quad\theta=S_{\tau}(\omega-{\textstyle\frac{1}{2}}\pi),$ (20a)
while for transpolar geodesics,
$\phi=\omega-{\textstyle\frac{1}{2}}\pi,\quad\tau={\textstyle\frac{1}{2}}\pi-\alpha,\quad\theta=S_{\tau}\beta.$ (20b)
We define
$f_{\theta}(\theta)$ $=\int_{0}\frac{\sqrt{1-\epsilon\kappa\cos^{2}\theta}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\theta}\sqrt{\kappa\cos^{2}\theta+\left|\mu\right|}}\,\mathrm{d}\theta,$ (21a)
$g_{\theta}(\theta)$ $=\int_{0}\frac{\kappa\cos^{2}\theta\sqrt{1-\epsilon\kappa\cos^{2}\theta}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\theta}\sqrt{\kappa\cos^{2}\theta+\left|\mu\right|}}\,\mathrm{d}\theta,$ (21b)
replacing $\phi$ by $\theta$ in Eqs. (18) and stipulating that positive square roots are to be taken in the integrands. We have replaced $\mu$ by $\left|\mu\right|$, consistent with the requirement that $0<\mu\leq\kappa^{\prime}$ for the $\theta$ integrals.
In the $\phi$ integrals in Eqs. (18), we have $\kappa\leq\mu<0$ leading to a weak (square-root) singularity in the integrand at $\cos\phi=\sqrt{\left|\mu\right|/\kappa}$. This singularity can be removed by changing the variable of integration to $\psi$ defined by
$\sin\phi$ $=\sin\psi\sqrt{1-\left|\mu\right|/\kappa},$ (22a)
$\frac{\mathrm{d}\phi}{\mathrm{d}\psi}$ $=\frac{\cos\psi}{\cos\phi}\sqrt{1-\left|\mu\right|/\kappa}.$ (22b)
Note that $\mathrm{d}\phi/\mathrm{d}\psi$ changes sign at the vertices of the geodesic, where $\cos\psi=0$ or, equivalently, where $\cos\tau=0$. It is convenient to write $\psi$ in terms of $\tau$ and $\theta$ by rewriting Eq. (14) as
$\mu=\kappa\cos^{2}\phi\sin^{2}\tau-\kappa^{\prime}\cos^{2}\theta\cos^{2}\tau.$ (23)
The interconversions of $\psi$, $\phi$, and $\tau$ become
[…]
The heavy ratio line in the argument to the arctangent indicates that the quadrant of the function is given by the signs of the numerator and denominator separately. This ensures that $\psi$ increases in the forward direction along a geodesic.
The functions $f(\phi)$ and $g(\phi)$ are replaced by
$f_{\psi}(\psi)$ $=\int_{0}\frac{\sqrt{1-\epsilon(\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi)}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi}}$
$\qquad\times\frac{\mathrm{d}\psi}{\sqrt{\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi}},$ (25a)
$g_{\psi}(\psi)$ $=\int_{0}\frac{\sqrt{1-\epsilon(\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi)}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi}}$
$\qquad\times\sqrt{\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi}\,\mathrm{d}\psi;$ (25b)
positive square roots should be taken in these integrals. The variable $\psi$ (which, like $\theta$, is a rotating coordinate) plays the same role as the arc length on the auxiliary sphere in Bessel’s solution of the geodesic problem on a biaxial ellipsoid; it allows the geodesic to be tracked through its vertices (points of extreme latitude) and to be followed indefinitely.
Now, the geodesic is given by
$\delta$ $=f_{\psi}(\psi)-f_{\theta}(\theta),$ (26a)
$(s+s_{1})/b$ $=g_{\psi}(\psi)+g_{\theta}(\theta).$ (26b)
The omitted parameters for $f_{\psi}$ and $g_{\psi}$ are $\kappa,\epsilon,\mu]=[k^{2},e^{2},-\gamma]$ for circumpolar geodesics and $k^{\prime 2},-e^{2},\gamma]$ for transpolar geodesics; conversely, the omitted parameters for $f_{\theta}$ and $g_{\theta}$ are $k^{2},e^{2},-\gamma]$ for transpolar geodesics and $k^{\prime 2},-e^{2},\gamma]$ for circumpolar geodesics.
The integrands in the definitions of these functions are analytic, positive, even, and periodic with period $\pi$; so the functions are analytic, odd, increasing functions, consisting of a secular term with a superimposed periodic ripple. (The integrand for $g_{\theta}$ vanishes for $\cos\theta=0$; so $g_{\theta}$ is merely non-decreasing.) We discuss the numerical evaluation of these integrals in Sec. 8.
It is clear that we have a complete solution to the direct problem for $\gamma\neq 0$. The initial conditions, $s=0$, $\beta_{1}$, $\omega_{1}$, $\alpha_{1}$, allow the constants $\delta$ and $s_{1}$ in Eqs. (26) to be determined. For a given distance from the initial point $s=s_{12}=s_{2}-s_{1}$, these equations have a unique solution for the endpoint $\theta_{2}$ and $\phi_{2}$, which allows the coordinates and azimuth at point 2 to be found.
6 Properties of umbilical geodesics
Umbilical geodesics are characterized by $\gamma=0$. In this case, the two integrals in Eq. (26a) have logarithmic singularities at $\cos\theta=0$ and $\cos\psi=0$. These singularities cancel at umbilics where both $\cos\theta$ and $\cos\psi$ vanish. A geodesic leaving a particular umbilic in a specified direction must arrive at the opposite umbilic with a well-defined azimuth. We use a connection relation to determine the azimuth on leaving that umbilic. The process can be repeated to follow a geodesic through multiple passages of the umbilics.
We treat the umbilical geodesics as the limiting case of circumpolar geodesics, i.e., $\gamma\rightarrow 0+$. (Treating the other case, $\gamma\rightarrow 0$-, follows a comparable procedure.) Thus, $\psi$ and $\theta$ are related to $\beta$ and $\omega$, respectively. With $\gamma=0$, both $(f_{\psi},g_{\psi})$ and $(f_{\theta},g_{\theta})$ have the same functional forms, with, for example, Eqs. (25) becoming
$f_{\psi}(\psi;\mu=0)$ $=\int_{0}\frac{\sqrt{1-\epsilon\kappa\cos^{2}\psi}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}\sqrt{\kappa}\cos\psi}\,\mathrm{d}\psi,$ (27a)
$g_{\psi}(\psi;\mu=0)$ $=\int_{0}\frac{\sqrt{\kappa}\cos\psi\sqrt{1-\epsilon\kappa\cos^{2}\psi}}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}}\,\mathrm{d}\psi.$ (27b)
We have substituted $\sqrt{\cos^{2}\psi}=\cos\psi$, because, within each geodesic segment, we will take $\left|\psi\right|\leq{\textstyle\frac{1}{2}}\pi$, so that $\cos\psi\geq 0$.
The integrals Eqs. (27) can be given in closed form in the spherical limit, $\epsilon=0$, giving
$f_{\psi}(\psi;\epsilon=0,\mu=0)$ $=\int_{0}\frac{1}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}\sqrt{\kappa}\cos\psi}\,\mathrm{d}\psi$
$=\frac{\sinh^{-1}(\sqrt{\kappa^{\prime}}\tan\psi)}{\sqrt{\kappa\kappa^{\prime}}},$ (28a)
$g_{\psi}(\psi;\epsilon=0,\mu=0)$ $=\int_{0}\frac{\sqrt{\kappa}\cos\psi}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}}\,\mathrm{d}\psi$
$=\tan^{-1}\frac{\sqrt{\kappa}\sin\psi}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}}.$ (28b)
We write
$f_{\psi}(\psi;\mu=0)$ $=f_{\psi}(\psi;\epsilon=0,\mu=0)-\Delta f_{\psi}(\psi;\mu=0),$ (29a)
$\Delta f_{\psi}(\psi;\mu=0)$ $=\int_{0}\frac{\epsilon\sqrt{\kappa}\cos\psi}{\sqrt{\kappa^{\prime}+\kappa\cos^{2}\psi}\bigl(1+\sqrt{1-\epsilon\kappa\cos^{2}\psi}\bigr)}\,\mathrm{d}\psi.$ (29b)
We could also express $g_{\psi}(\psi;\mu=0)$ as the sum of $g_{\psi}(\psi;\allowbreak{\epsilon=0},{\mu=0})$ and a correction, but this is not necessary for its accurate evaluation. The integrand in Eq. (29b) is free of singularities, so Eq. (28) captures the full singular behavior of $f_{\psi}(\psi;{\mu=0})$. In the limit $\psi\rightarrow\pm{\textstyle\frac{1}{2}}\pi$, we have
$f_{\psi}(\psi;\mu=0)\rightarrow\pm\frac{\log(2\sqrt{\kappa^{\prime}}\sec\psi)}{\sqrt{\kappa\kappa^{\prime}}}.$ (30)
For each of Eqs. (27)–(30), we have a corresponding equation substituting $\theta$ for $\psi$. In the functions of $\psi$, the implied parameters are $\kappa,\epsilon]=[k^{2},e^{2}]$, and for the functions of $\theta$, the parameters are $k^{\prime 2},-e^{2}]$.
Label each geodesic segment from one umbilic to the next sequentially with index $j$. We use superscripts $\mp$ to label the start ($\theta=\psi=-{\textstyle\frac{1}{2}}\pi$) and end ($\theta=\psi={\textstyle\frac{1}{2}}\pi$) of a segment. At either end of the $j$th segment, we have
$\delta_{j}$ $=\pm\biggl(\frac{1}{kk^{\prime}}\log\frac{k^{\prime}\sec\psi_{j}^{\pm}}{k\sec\theta_{j}^{\pm}}-\Delta f_{\psi}({\textstyle\frac{1}{2}}\pi)+\Delta f_{\theta}({\textstyle\frac{1}{2}}\pi)\biggr)$
$=\pm\frac{1}{kk^{\prime}}\biggl(\log\left|\tan\alpha_{j}^{\pm}\right|-{\textstyle\frac{1}{2}}\Delta\biggr),$ (31)
where
$\Delta=2kk^{\prime}\bigl(\Delta f_{\psi}({\textstyle\frac{1}{2}}\pi)-\Delta f_{\theta}({\textstyle\frac{1}{2}}\pi)\bigr).$ (32)
We have introduced $\alpha_{j}^{\pm}$ from Eq. (11), which gives
$\frac{k^{\prime}\sec\psi}{k\sec\theta}=\pm\tan\alpha.$ (33)
Equating $\delta_{j}$ at the two ends of the umbilical segment gives
$\tan\alpha_{j}^{-}\tan\alpha_{j}^{+}=\exp(\Delta),$ (34)
where $\alpha_{j}^{\pm}$ lie in the same quadrant.
To connect to the next segment, we need to jump over the umbilic. Consider the point near the end of a geodesic segment at $(\beta_{j}^{+},\omega_{j}^{+})$. The point at the start of the next segment at $(\beta_{j+1}^{-},\omega_{j+1}^{-})$ is on the diametrically opposite side of the umbilic and is given by
$\cos\beta_{j+1}^{-}=\pm\frac{k^{\prime}}{k}\sin\omega_{j}^{+},\qquad\sin\omega_{j+1}^{-}=\mp\frac{k}{k^{\prime}}\cos\beta_{j}^{+},$ (35)
with the sign of $\cos\beta$ preserved. The azimuths are related by
$\tan\alpha_{j}^{+}\tan\alpha_{j+1}^{-}=-1,\qquad\alpha_{j+1}^{-}-\alpha_{j}^{+}=\pm{\textstyle\frac{1}{2}}\pi,$ (36)
with the sign chosen to preserve the sign of $\sin\alpha$.
This allows us to express $\delta_{j+1}$, using the lower signs in Eq. (31), in terms of $\delta_{j}$, using the upper signs in Eq. (31), which gives
$\delta_{j+1}=\delta_{j}+\frac{\Delta}{kk^{\prime}}.$ (37)
In general, we obtain
$\delta_{j}$ $=\delta_{0}+j\frac{\Delta}{kk^{\prime}},$ (38a)
$\tan\alpha_{j}^{-}$ $=(-1)^{j}\exp(-j\Delta)\tan\alpha_{0}^{-},$ (38b)
$\tan\alpha_{j}^{+}$ $=(-1)^{j}\exp\bigl((j+1)\Delta\bigr)/\tan\alpha_{0}^{-}.$ (38c)
The presence of the exponential terms in Eqs. (38) is evidence of the instability of umbilical geodesics discussed in Sec. 3. This property of umbilical geodesics was discovered by Hart (1849), who gives an alternative (but equivalent) expression for $\Delta$,
$\Delta=\int_{0}^{\pi/2}\frac{\sqrt{(a^{2}-b^{2})(b^{2}-c^{2})}}{\sqrt{a^{2}\tan^{2}\phi+b^{2}}\sqrt{c^{2}\tan^{2}\phi+b^{2}}}\,\mathrm{d}\phi.$ (39)
The integrands in Eqs. (27b) and (29b) are analytic, even, and periodic with period $2\pi$. In addition, they are odd about the point $\theta={\textstyle\frac{1}{2}}\pi$ and positive for $\left|\theta\right|<{\textstyle\frac{1}{2}}\pi$. Thus, the integrals are analytic, odd, periodic with period $2\pi$, and increasing in the interval $\left|\theta\right|<{\textstyle\frac{1}{2}}\pi$. The distance between opposite umbilics is
$s_{0}=2b\bigl(g_{\psi}({\textstyle\frac{1}{2}}\pi;\mu=0)+g_{\theta}({\textstyle\frac{1}{2}}\pi;\mu=0)\bigr),$ (40)
half the perimeter of the median principal ellipse.
When solving the direct problem, $\delta$ and $s_{1}$ are determined as before, with the proviso that, if the initial point is an umbilic, $\delta=\delta_{0}$ should be evaluated using Eq. (31) and the initial azimuth. With a given $s_{12}$, determine $s_{2}=s_{1}+s_{12}$ and find the umbilical segment index with
$j=\lfloor s_{2}/s_{0}\rceil,$ (41)
where $\lfloor x\rceil$ is the nearest integer to $x$. Then solve the geodesic equations with $s=s_{2}-js_{0}$ and $\delta=\delta_{j}$ found with Eq. (38a). If initial conditions are such that the geodesic lies on the median principal ellipse, then $\delta=\pm\infty$ diverges. In this case, the geodesic can be broken into segments of lengths $2bg_{\psi}({\textstyle\frac{1}{2}}\pi;{\mu=0})$ and $2bg_{\theta}({\textstyle\frac{1}{2}}\pi;{\mu=0})$, the distances between neighboring umbilics.
7 Biaxial ellipsoids
Geodesics on a biaxial ellipsoid are well understood. However, it’s instructive to see how we can recover the biaxial solution from Jacobi’s.
For biaxial ellipsoids, the functions $f_{\theta}$ and $g_{\theta}$ are evaluated with $\kappa=0$ to give
$f_{\theta}(\theta;\kappa=0)$ $=\frac{\theta}{\sqrt{\left|\gamma\right|}},$ (42a)
$g_{\theta}(\theta;\kappa=0)$ $=0.$ (42b)
The functions $f_{\psi}$ and $g_{\psi}$, with $\kappa=1$, become
$f_{\psi}(\psi;\kappa=1)$ $=\int_{0}\frac{\sqrt{1-\epsilon(\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi)}}{\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi}\,\mathrm{d}\psi,$ (43a)
$g_{\psi}(\psi;\kappa=1)$ $=\int_{0}\sqrt{1-\epsilon(\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi)}\,\mathrm{d}\psi.$ (43b)
In the spherical limit, we carry out the integration to give
$f_{\psi}(\psi;\kappa=1,\epsilon=0)$ $=\int_{0}\frac{1}{\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi}\,\mathrm{d}\psi$
[…]
$g_{\psi}(\psi;\kappa=1,\epsilon=0)$ $=\psi.$ (44b)
(The arctangent function in Eq. (44a) tracks the quadrant of $\psi$ through multiple revolutions.) Following the same procedure used in the umbilical case, Eqs. (29), we write
$f_{\psi}(\psi;\kappa=1)$ $=f_{\psi}(\psi;\kappa=1,\epsilon=0)-\Delta f_{\psi}(\psi;\kappa=1),$ (45a)
$\Delta f_{\psi}(\psi;\kappa=1)$ $=\int_{0}\frac{\epsilon}{1+\sqrt{1-\epsilon(\cos^{2}\psi\left|\gamma\right|\sin^{2}\psi)}}\,\mathrm{d}\psi.$ (45b)
The geodesic equations Eqs. (26) become
[…]
$\quad{}-\int_{0}\frac{\epsilon\sin\tau_{0}}{1+\sqrt{(1-\epsilon)+\epsilon\cos^{2}\tau_{0}\sin^{2}\psi}},$ (46a)
$s+s_{1}$ $=b\int_{0}\sqrt{(1-\epsilon)+\epsilon\cos^{2}\tau_{0}\sin^{2}\psi}\,\mathrm{d}\psi,$ (46b)
where $\sin^{2}\tau_{0}=\left|\gamma\right|=\cos^{2}\phi\sin^{2}\tau$, $\epsilon=e^{2}=({a^{2}-c^{2}})/a^{2}$ for oblate ellipsoids, and $\epsilon=-e^{2}=({c^{2}-a^{2}})/c^{2}$ for prolate ellipsoids. It is readily seen that these agree with the standard formulas for biaxial ellipsoids, e.g., Eqs. (8) and (7) of Karney (2013).
We treat meridional geodesics on a biaxial ellipsoid by the same mechanisms used for umbilical geodesics in Sec. 6. In the limit $\gamma\rightarrow 0$, Eq. (44a) becomes
$f_{\psi}(\psi;\kappa=1,\epsilon=0,\mu\rightarrow 0)=\frac{\lfloor\psi/\pi\rceil\pi}{\sqrt{\left|\gamma\right|}}+\tan\psi.$ (47)
Equation (26a) gives the expected result that $\theta$ is constant except on passage through a pole, $\cos\psi=0$, where $\theta$ increases by $\pi$. Passages through the poles are at intervals of $s_{0}=2bg_{\psi}({{\textstyle\frac{1}{2}}\pi;{\kappa=1},{\mu=0}})$. The solution of the direct problem proceeds analogously to the umbilical case.
When tackling the inverse problem, we will need to find meridional conjugate points. (This is needed to determine whether the shortest geodesic for two points on opposite meridians follows the meridian.) The conjugate points correspond to a change in $\theta$ by a multiple of $\pi$, giving
$f_{\theta}(\theta_{2};\kappa=0)-f_{\theta}(\theta_{1};\kappa=0)=\frac{n\pi}{\sqrt{\left|\gamma\right|}}.$ (48)
To cancel this singular change in $f_{\theta}$, $\psi$ must cross over $n$ poles. Assuming that $\psi_{1}\in(-{\textstyle\frac{1}{2}}\pi,{\textstyle\frac{1}{2}}\pi)$, we have $\psi_{2}\in\bigl(({n-{\textstyle\frac{1}{2}}})\pi,\allowbreak({n+{\textstyle\frac{1}{2}}})\pi\bigr)$. To find the value of $\psi$ within the allowed range, we balance the second non-singular term in Eq. (47) against $\Delta f_{\psi}$, i.e., we solve
$\tan\psi_{2}-\tan\psi_{1}$ $=\Delta f_{\psi}(\psi_{2};\kappa=1,\mu=0)$
$\qquad{}-\Delta f_{\psi}(\psi_{1};\kappa=1,\mu=0),$ (49)
for $\psi_{2}$ (in the allowed range). This is accomplished by solving for $y=\tan\psi_{2}$ using Newton’s method.
8 Evaluating the integrals
Jacobi’s solution reduces the original geodesic problem, coupled ordinary differential equations, to the evaluation of one-dimensional integrals. Jacobi identifies these as abelian integrals, but this does not particularly help because there are no simple procedures for computing them.
We are therefore left with numerical quadrature of some sort. This is exactly the approach taken by Cayley (1872) to construct graphically the paths of umbilical geodesics (as shown in Fig. 3). More recently, Baillard (2015) provides routines for the HP-41 calculator for solving the inverse geodesic problem by performing the corresponding definite integrals for Jacobi’s solution using Gauss-Legendre quadrature.
In this work, I sought a method for evaluating the integrals that allows the solution to be found with nearly full double-precision accuracy. We would also like to be able to compute the indefinite integrals rapidly, so that points at arbitrary positions on a geodesic can be found. The method should lend itself to implementation at higher precision with a corresponding increase in accuracy at a modest cost.
The concept of indefinite integration in numerical applications was introduced by Clenshaw and Curtis (1960), who showed that, having computed a definite integral over the range $[a,b]$, it is possible with “little extra complication” to determine the integral over any interior interval. Their method naturally extends to periodic functions approximated by a Fourier series. The steps are: approximate the integrand with a Fourier series, with the coefficients found using the fast Fourier transform; trivially integrate the series; evaluate the integral at arbitrary points using Clenshaw summation (Clenshaw, 1955). Trefethen and Weideman (2014) review the mathematical background for why this method gives such high accuracy.
In some cases, the integrands are almost singular, e.g., the term $\sqrt{\kappa\cos^{2}\theta+\left|\mu\right|}$ in the denominator of Eq. (21a) leads to a sharp peak in the integrand when $\mu$ is very small, which in turn requires the inclusion of many terms in the Fourier series. This problem can be avoided by a change of variables using
$x$ $=F(\phi,q),$ $\phi$ $=\operatorname{am}(x,q),$ (50a)
$\frac{\mathrm{d}x}{\mathrm{d}\phi}$ $=\frac{1}{\sqrt{1-q^{2}\sin^{2}\phi}},$ $\frac{\mathrm{d}\phi}{\mathrm{d}x}$ $=\operatorname{dn}(x,q),$ (50b)
where $F(\phi,q)$ is the elliptic integral of the first kind, $\operatorname{am}(x,q)$ is the Jacobi amplitude function, and $\operatorname{dn}(x,q)$ and $\operatorname{cn}(x,q)$ (used below) are Jacobi elliptic functions. We adopt the notation of (Olver et al., 2010, Chaps. 19 & 22) except that, to avoid confusion with $k$ defined in Eq. (2), we use $q$ to denote the modulus.
Substituting
$\theta=\operatorname{am}\bigl(v,\sqrt{\kappa/(\kappa+\left|\mu\right|)}\bigr)$ (51)
in Eq. (21a) gives
$f_{\theta}(\theta)=\int_{0}^{\operatorname{am}\theta}\frac{\sqrt{1-\epsilon\kappa\operatorname{cn}^{2}v}}{\sqrt{\kappa^{\prime}+\kappa\operatorname{cn}^{2}v}\sqrt{\kappa+\left|\mu\right|}}\,\mathrm{d}v,$ (52)
where, for brevity’s sake, we have omitted the modulus $q=\kappa/({\kappa+\left|\mu\right|})$. The same change of variables is made with Eq. (21b).
In Eqs. (25), we substitute
$\psi=\operatorname{am}\bigl(u,\sqrt{(\kappa-\left|\mu\right|)/\kappa}\bigr),$ (53)
for $\left|\mu\right|$ small, to cancel the factor $\sqrt{\kappa\cos^{2}\psi+\left|\mu\right|\sin^{2}\psi}$ in Eqs. (25). Finally, in Eqs. (27b) and (29b), we substitute $\theta=\operatorname{am}(w,\sqrt{\kappa})$ for $\kappa^{\prime}$ small, to cancel the factor $\sqrt{\kappa^{\prime}+\kappa\cos^{2}\theta}$ in Eq. (29b).
Even though the cost of computing $F(\phi,q)$ and $\operatorname{am}(x,q)$ is small, we need only incur the cost when $q$ is sufficiently close to 1, e.g., $q^{2}>7/8$.
There is another instance where the integrands are nearly singular, namely for an almost flat ellipsoid with $c/b$ small, i.e., $ek$ close to unity. In this case, the factor $\sqrt{1-e^{2}k^{2}\cos^{2}\beta}$ appearing in Eqs. (17) has a sharp dip at $\beta=0$. The dip could be smoothed out by a suitable change of the variable of integration, but that will take us too far afield. The case where $c$ vanishes is discussed in Appendix D
9 The direct problem
The direct problem, determining the position at a given distance along a geodesic, is found by solving the nonlinear simultaneous equations Eqs. (26). This is accomplished using Newton’s method in two dimensions; details are given in Appendix C. We need, first, to specify the two independent variables to use.
For general geodesics $\gamma\neq 0$, the domains of $\psi$ and $\theta$ are unbounded. It is preferable to use the new integration variables, $u$ and $v$ defined by Eqs. (53) and (51), as the independent variables instead of $\psi$ and $\theta$ (assuming a change of variable was needed), because this results in smoother functions.
For umbilical geodesics, it is important to “transform away” the singular behavior of $f_{\psi}$ and $f_{\theta}$ at the umbilics by using
$u=\sinh^{-1}(k^{\prime}\tan\psi),\qquad v=\sinh^{-1}(k\tan\theta),$ (54)
as the independent variables, so that the leading spherical contributions Eqs. (28) become
$f_{\psi}(\psi;\epsilon=0,\mu=0)$ $=\frac{u}{kk^{\prime}},$ (55a)
$g_{\psi}(\psi;\epsilon=0,\mu=0)$ $=\tan^{-1}\frac{k\tanh u}{k^{\prime}}.$ (55b)
In solving problems with umbilical geodesics, we keep track of transits through umbilics so that we can make the restrictions, $\left|\psi\right|\leq{\textstyle\frac{1}{2}}\pi$ and $\left|\theta\right|\leq{\textstyle\frac{1}{2}}\pi$. The problem now maps to the infinite domain in $(u,v)$ coordinates, as is the case for general geodesics.
There are a few special cases in dealing with umbilical geodesics: the geodesic lies on the median ellipse for which $\delta\rightarrow\infty$ or the target point is an umbilic, in which case $u$ and $v$ diverge. The strategy in these cases is to let $\delta$, $u$, or $v$ take on a large but finite value, so that, for example, inverting Eq. (54) gives $\psi=\pm{\textstyle\frac{1}{2}}\pi$ with high accuracy, but not so large that $\sinh u$ overflows. I find that values close to $3\log\epsilon$ suffice; here $\epsilon$ is the machine epsilon, typically $2^{-52}$. In this case, $\cos\psi=k^{\prime}/\sqrt{\sinh^{2}u+k^{\prime 2}}$ is nonzero (and similarly for $\cos\theta$), allowing $\alpha$ to be determined from Eq. (11).
Turning to the special case of biaxial ellipsoids, we treat them, as much as possible, the same as triaxial ellipsoids. This means a slightly different treatment for meridional geodesics compared with umbilical geodesics. One other change was found to help in maintaining accuracy. Normally, $f_{\psi}(\psi;{\kappa=1})$ is computed by substituting $\kappa=1$ into Eq. (25a). For triaxial ellipsoids with small $\gamma$, we make the change of variables Eq. (53) to smooth out the integrand. In the corresponding situation for biaxial ellipsoids, it’s better to determine $f_{\psi}(\psi;{\kappa=1})$ using Eqs. (45) and (44a), where the near-singular term is subtracted from the integrand and integrated analytically.
10 The inverse problem
[…]
[scale=0.75]figures/cuta
[…]
[scale=0.75]figures/cutb
Figure 5: Geodesics emanating from a single point on an ellipsoid. The ellipsoid and viewpoint are the same as Fig. 1. The geodesics start at $\beta_{1}=-34.46^{\circ}$, $\omega_{1}=-149.94^{\circ}$, and the azimuths $\alpha_{1}$ are multiples of $7.5^{\circ}$. The geodetic coordinates of the starting point are $\phi_{1}=-40^{\circ}$, $\lambda_{1}=30^{\circ}-180^{\circ}$, the opposite of the viewing direction. Part (a) shows the geodesics followed up to the points where they are no longer the shortest geodesics. The union of such points, the cut locus, is shown as a heavy line and is a segment of the line of curvature $\beta=-\beta_{1}$. Part (b) shows the geodesics from (a) as dotted lines, and these are continued (as solid lines) until they meet at the line of curvature $\omega=\omega_{1}+180^{\circ}$ (shown as a heavy line).
[…]
[scale=0.75]figures/cutunroll
[…]
[scale=0.75]figures/hybrid
Figure 6: (a) The geodesics in Fig. 5(a) displayed in a plate carrée (longitude-latitude) projection. The geodesics all start at $(\beta_{1},\omega_{1})$, with azimuths $\alpha_{1}$ which are multiples of $7.5^{\circ}$, and are continued until they encounter the cut locus (shown as a heavy line). For a given $\beta_{2}\in[-\left|\beta_{1}\right|,\left|\beta_{1}\right|]$, the longitude $\omega_{2}$ is an increasing function of $\alpha_{1}$. (b) The mapping from $\alpha_{1}$ to $\omega_{2}$ for the geodesics in (a). The starting coordinates $(\beta_{1},\omega_{1})$ are fixed, and curves of $\omega_{2}=\omega_{2}^{*}(\alpha_{1};{\beta_{1},\omega_{1},\beta_{2}})$ are shown for $\beta_{2}=[-1,-0.6,0,0.6,1]\times\left|\beta_{1}\right|$. The corresponding plot for a biaxial ellipsoid is given by Karney (2013, Fig. 4).
Crucial to solving the inverse problem is an understanding of the properties of all the geodesics emanating from a single point $(\beta_{1},\omega_{1})$.
  • If the starting point is an umbilic, all the lines meet at the opposite umbilic at a distance $s_{0}$ given by Eq. (40).
  • Otherwise, the first envelope of the geodesics is a four-pointed astroid; see Fig. 5(b). Two of the cusps of the astroid lie on $\beta=-\beta_{1}$, and the other two lie on $\omega=\omega_{1}+\pi$. This is the so-called “last geometric statement” of Jacobi (1843, §6).
  • All the geodesics intersect (or, in the case of $\alpha_{1}=0$ or $\pi$, touch) the line $\omega=\omega_{1}+\pi$.
  • All the geodesics intersect (or, in the case of $\alpha_{1}=\pm{\textstyle\frac{1}{2}}\pi$, touch) the line $\beta=-\beta_{1}$.
  • The two geodesics with azimuths $\pm\alpha_{1}$ first intersect on $\omega=\omega_{1}+\pi$, and their lengths to the point of intersection are equal.
  • The two geodesics with azimuths $\alpha_{1}$ and $\pi-\alpha_{1}$ first intersect on $\beta=-\beta_{1}$, and their lengths to the point of intersection are equal.
The last property defines the cut locus for $(\beta_{1},\omega_{1})$; this is the locus of points where the geodesics cease to be shortest geodesics. This is a segment of a line of curvature $\beta=-\beta_{1}$; see Fig. 5(a). This figure shows the shortest geodesic between $(\beta_{1},\omega_{1})$ and any other point $(\beta_{2},\omega_{2})$ on the ellipsoid. Without loss of generality, we take $\beta_{1}\leq 0$. Then, for $\beta_{2}\in[-\left|\beta_{1}\right|,\left|\beta_{1}\right|]$, each geodesic intersects the line $\beta=\beta_{2}$ exactly once. For a given $\beta_{2}$, this defines a monotonic mapping $\omega_{2}^{*}(\alpha_{1};{\beta_{1},\omega_{1},\beta_{2}})$ of the circle of azimuths $\alpha_{1}$ to the circle of longitudes $\omega_{2}$; see Fig. 6. The mapping is continuous except if $\beta_{1}=0$- and $\beta_{2}=0$, where, for example, $\omega_{2}^{*}$ jumps from $\omega_{1}$, for $\alpha_{1}={\textstyle\frac{1}{2}}\pi$-, to the conjugate longitude, for $\alpha_{1}={\textstyle\frac{1}{2}}\pi+$.
These properties show that the inverse problem can be solved using techniques almost the same as those employed for a biaxial ellipsoid. The first task is to treat those cases where both endpoints are on one of the principal ellipses to determine whether the shortest geodesic lies on that ellipse.
Starting with the median principal ellipse, we have the following possibilities:
  • If the points are opposite umbilics, an arbitrary $\alpha_{1}$ may be chosen. Two of the shortest paths follow the median ellipse. However, it is more useful to pick the geodesic which contains the point $\beta=0$, $\omega={\textstyle\frac{1}{2}}\pi$; this gives finite and nonzero values for $\tan\alpha_{1}$ and $\tan\alpha_{2}$, allowing other shortest geodesics to be generated.
  • Otherwise, if the shortest path on the ellipse crosses no more than a single umbilic, the shortest geodesic lies on the median ellipse. This includes all cases where at least one of the endpoints is an umbilic.
  • Otherwise, if the endpoints satisfy $\sin\omega_{1,2}=0$ (the points are near $X=\pm a$), the shortest geodesic is on the median ellipse.
  • Otherwise, if the endpoints satisfy $\cos\beta_{1,2}=0$ (the points are near $Z=\pm c$), the shortest geodesic is on the median ellipse only if there is no intervening conjugate point. Otherwise, there are two shortest paths, and one of them can be found using the general method (given below) with the azimuth restricted to, e.g., $\alpha_{1}\in(-{\textstyle\frac{1}{2}}\pi,{\textstyle\frac{1}{2}}\pi)$, assuming that $\beta_{1}=-{\textstyle\frac{1}{2}}\pi)$.
We turn now to the other principal ellipses. The case where both points are on the major ellipse, $\beta_{1,2}=0$, is treated in the same way as the last subcase for the median ellipse. The shortest path is along the ellipse, provided there is no intervening conjugate point. Otherwise, there are two shortest paths, one of which has e.g., $\alpha_{1}\in(-{\textstyle\frac{1}{2}}\pi,{\textstyle\frac{1}{2}}\pi)$, and this is found using the general method. If both points are on the minor principal ellipse, the shortest path always follows the ellipse.
There is another instance where the azimuths can be found directly, namely, if only one of the endpoints is an umbilic. We then have $\gamma=0$, and the azimuth at the other point is given by Eq. (11), picking the signs of the sine and cosine appropriately.
We now come to the general case where the shortest geodesic does not lie on a principal ellipse and neither point is an umbilic. The process closely follows Karney (2013, §4). Using symmetries, arrange that $\beta_{1}\leq 0$ and $\left|\beta_{1}\right|\leq\beta_{2}\leq\left|\beta_{1}\right|$, the situation depicted in Fig. 6. Find the azimuth $\alpha_{1}$ at point 1 which satisfies
$\omega_{2}^{*}(\alpha_{1};\beta_{1},\omega_{1},\beta_{2})=\omega_{2}\pmod{2\pi}$; (56)
this is a one-dimensional root-finding problem. Finally, determine the length of the geodesic segment.
A few remarks are in order:
  • Finding the root $\alpha_{1}$ of Eq. (56), which may require several iterations, only requires consideration of Eq. (26a), which determines the course of the geodesic. The calculation of the distance using Eq. (26b) can be postponed until $\alpha_{1}$ has been found.
  • With the biaxial problem, Helmert (1880, Eq. (6.5.1)) provided a formula for the reduced length $m_{12}$, allowing conjugate points to be found by the condition, $m_{12}=0$. I know of no corresponding formula for $m_{12}$ for the triaxial case. Nevertheless, it is possible to find the first conjugate point by the condition that $\beta$ has completed half an oscillation, which in turn implies that $\psi$ has advanced by $\pi$. Equation (26a) can be used to give the value of $\omega$ at that point. In the special case of a biaxial ellipsoid, the conjugate point for a meridional geodesic is given by Eq. (49); this is only needed for solving inverse problems on a prolate ellipsoid.
  • In the biaxial case, we were able to solve Eq. (56) using Newton’s method because the necessary derivative was given in terms of $m_{12}$. In the triaxial case, we do not have an expression for $m_{12}$, so we resort to a simpler root-finding method.
  • For biaxial ellipsoids, rotate the points about the axis of symmetry so that $\omega_{1}=0$ (resp. $\beta_{1}=-{\textstyle\frac{1}{2}}\pi$) for oblate (resp. prolate) ellipsoids. In this case, meridional geodesics are handled by the logic for the median principal ellipse.
  • The treatment of prolate ellipsoids differs from previous work (Karney, 2013). There, the generalization of the oblate case entailed finding the intersection of a geodesic with a circle of geodetic latitude, which corresponds to a circle of ellipsoidal longitude. In this triaxial treatment, we find the intersection with a line of constant ellipsoidal latitude, which corresponds to a meridian on a prolate ellipsoid.
The shortest path is unique unless:
  • The length of the geodesic vanishes $s_{12}=0$, in which case any constant can be added to the azimuths.
  • The points are opposite umbilics (this only applies for triaxial ellipsoids, i.e., $k\neq 0$ and $k^{\prime}\neq 0$). In this case, $\alpha_{1}$ can take on any value, and $\alpha_{2}$ needs to be adjusted to maintain the value of $\tan\alpha_{1}/\tan\alpha_{2}$.
  • $\beta_{1}+\beta_{2}=0$ and $\cos\alpha_{1}$ and $\cos\alpha_{2}$ have opposite signs. In this case, there is another shortest geodesic with azimuths $\pi-\alpha_{1}$ and $\pi-\alpha_{2}$.
Any azimuth can be used for the shortest path connecting two opposite poles on a biaxial ellipsoid or any two opposite points on a sphere.
There is an additional interesting property of geodesics: the geodesic distance between the points $(\beta_{1},\omega_{1})$ and $(\beta_{2},\omega_{2})$ equals that between $(\beta_{2},\omega_{1})$ and $(\beta_{1},\omega_{2})$. This is a consequence of evaluating the integrals appearing in Eqs. (17) between the same limits in the two cases. We can state this another way: Consider a curvilinear rectangle whose sides are lines of curvature; the diagonals of this rectangle are equal. This is known as Ivory’s Lemma.
11 Implementation
An implementation of the solutions of the direct and inverse geodesic problems is provided with version 2.6 of the C++ library, GeographicLib (Karney, 2025). It’s practically impossible to exhaustively document the methods here; the reader is referred to the code for details. Here, I give an overview of the important aspects of the code.
The class for performing the Fourier approximations to the integrands in Jacobi’s solution was inspired by the support for periodic functions that was added to Chebfun (Wright et al., 2015). The number of sample points is successively doubled until convergence as defined by the “chopping” criterion given by Aurentz and Trefethen (2017). Constructing a Fourier series for the integral is simple.
I provide an optimized computation of the inverse of the integral. This uses Newton’s method to compute a single value of the inverse. During the course of refining the Fourier series, a significant speedup is achieved by using the current Fourier series for the inverse to provide accurate starting guesses of the inverse at the new sample points. The last (and most costly) rounds of Fourier refinement require only a single Newton iteration for each sample point. My initial expectation was that this would be useful in cases where many waypoints needed to be found. In the event, I used the two-dimensional Newton’s method as described in Appendix C to solve the direct problem; this obviates the need for finding the Fourier series for the inverse.
Turning to the solution of the inverse problem, I will focus on the general case. As described in Sec. 10, this involves solving Eq. (56) for $\alpha_{1}$. The solution of Eq. (56) starts by finding $\omega_{2}^{*}$ for the four umbilical directions (these all use the same $f_{\psi}$ and $f_{\theta}$), and these serve to bracket the solution $\alpha_{1}$. The root is found by the method given by Chandrupatla (1997).
The solutions to the direct and inverse problems also return a “geodesic line” object. This holds the four functions $f_{\psi}$, $f_{\theta}$, $g_{\psi}$, and $g_{\theta}$, and the constants $\delta$ and $s_{1}$; this allows waypoints along the geodesic to be computed efficiently.
Testing for geodesics on a triaxial ellipsoid is about two orders of magnitude more challenging than the biaxial case. The shape of the ellipsoid is specified by 2 parameters ($e$ and $k$) instead of just one (the flattening), and the solution of the inverse problem depends on the longitudes of both endpoints instead of just their difference. So I limited my initial testing to a single ellipsoid, Cayley’s ellipsoid scaled to the median semiaxis, i.e., $a,b,c]=[\sqrt{2},1,1/\sqrt{2}]$, with an emphasis on exploring all the different inverse problems outlined in Sec. 10. The test set (Karney, 2024c) contains 500 000500\,000 geodesics, computed at high precision, with the coordinates of the endpoints given as integer degrees. My testing also included other ellipsoids with $a/c=2$, including oblate and prolate ellipsoids. In addition, I tested with spheres ($a=c$) with different values of $k$, and with a triaxial model of the earth (Panou et al., 2020). In the following, I will only report results for the published dataset for Cayley’s ellipsoid. These should be regarded as rough indications of those likely to be obtained with ellipsoids which are not too eccentric.
Averaging over the data in the test set, the mean number of Fourier coefficients required to represent any of the functions $f_{\psi}$, $f_{\theta}$, $g_{\psi}$, or $g_{\theta}$ is about 30. The solution of the two-dimensional equations for the direct geodesic problem using Newton’s method requires, on average, 5 iterations. The solution of the inverse method requires an average of 8 iterations of Chandrupatla’s method.
These figures are for double precision with the error tolerance set to machine precision, so that nearly full double-precision accuracy is achieved. Repeating the tests at higher precisions (64, 113, and 256 bits of precision, instead of 53 for double precision) shows that the number of Fourier coefficients scales proportionally to the number of bits. Two-dimensional Newton’s method for the direct problem enjoys the expected quadratic convergence: the error is squared on each iteration. The convergence is somewhat slower for the iterative solution for the inverse problem; on each iteration, the logarithm of the error is multiplied by about 7/4 (versus 2 for quadratic convergence).
When assessing the errors, it makes little sense to directly compare the ellipsoidal coordinates and azimuths because these vary very rapidly near the umbilics. A better approach is to compare these quantities expressed as cartesian positions and directions. The differences are given in “units in the last place” ($\mathrm{ulp}$), which I define to be $b/2^{53}$ for the error in the position (this is also used for the errors in the distance returned by the solution of the inverse problem) and $1/2^{53}\,\mathrm{rad}$ for the errors in the direction. For the earth, $1\,\mathrm{ulp}$ corresponds to about $0.7\,\mathrm{nm}$. The errors, so quoted, will be approximately the same as for long double precision (with $2^{53}$ replaced by $2^{64}$ to match the increase in precision).
For the test set for Cayley’s ellipsoid, the mean error in the position and direction returned for the solution to the direct problem is $5\,\mathrm{ulp}$ and $6\,\mathrm{ulp}$, respectively. For the solution to the inverse problem, the mean error in the distance is $3\,\mathrm{ulp}$. For this case, we do not compare the azimuths to the test data, because, for example, there may be multiple allowed azimuths. Instead, we demand consistency in the forward and backward direct problems given by the inverse solution, measuring the discrepancies in the positions and directions at the opposite endpoint; these are $6\,\mathrm{ulp}$ for the positions and $7\,\mathrm{ulp}$ for the directions.
These mean errors are impressively small. However, for practical applications, we need to quantify the maximum errors. Using the test data can only give a lower bound because the data offers rather sparse coverage; in particular, it omits potentially problematic geodesics that pass very close to umbilics. Based just on the test set, the maximum errors in the position and direction for the direct problem are $160\,\mathrm{ulp}$ and $1500\,\mathrm{ulp}$. The maximum error in the distance for the inverse problem is $90\,\mathrm{ulp}$; the maximum discrepancy in the position with the resulting direct problems is $9000\,\mathrm{ulp}$.
These maximum errors are still reasonably small. But more exhaustive testing on this and other ellipsoids will surely uncover instances where the errors are larger. A reasonable course would be to assess the errors in the context of a specific application. This will have the advantage of narrowing the parameters for the ellipsoids and will set a definite limit on the acceptable errors. If the errors are too large, the present implementation can be used at higher precision to help track down where the errors are creeping in.
[…]
An implementation of the coordinate conversions given in Appendix A is included in GeographicLib, as is sample code for solving the ordinary differential equations for the direct geodesic problem (Appendix E).
12 Discussion
In this paper, I have described the implementation of Jacobi’s solution to the direct geodesic problem for a triaxial ellipsoid. The method involves using Fourier series to represent the integrands appearing in Jacobi’s solution, which allows the integrals to be evaluated very accurately. In addition, I show how the bisection method can be applied to Newton’s method in two dimensions, which allows the coupled system of nonlinear equations for the direct problem to be solved efficiently. The solution for the inverse problem follows the same basic method as for biaxial ellipsoids.
The code is only about ten times slower than the much simpler case of the biaxial ellipsoid. I had also hoped to be able to say that the code is only ten times less accurate than the biaxial case. While the average errors do meet this condition, there are cases where the errors are substantially larger; this requires more study. Nevertheless, a key goal has been met: if necessary, the algorithms can be run with high-precision arithmetic, at a reasonable cost, to obtain more accurate results.
An alternative method for solving geodesics is to integrate the ordinary differential equations (ODEs) directly. This is most easily carried out in cartesian coordinates as advocated by Panou and Korakitis (2019). Some data on using this approach are given in Appendix E. The distinctions are as follows:
  • The code for solving the ODEs is considerably simpler, provided that a good “off-the-shelf” library for ODEs is available.
  • The ODEs only provide a solution for the direct geodesic problem. It is possible to extend this method to solve the inverse problem, but this increases the complexity considerably.
  • Jacobi’s solution is somewhat more accurate for typical distances. For long geodesics, the accuracy of Jacobi’s solution is maintained, while the ODE solution progressively degrades.
  • The ODE solution solves the direct problem somewhat faster than the Jacobi solution for typical distances. However, the CPU time for Jacobi’s solution is independent of distance, while it is proportional to distance for the ODE solution.
  • The ODE solution can be easily extended to compute the reduced length and the geodesic scale; see Appendix F.
Appendix A Coordinates for points on the ellipsoid
We consider the coordinates for points on the surface of the ellipsoid. (Extending the coordinate system to treat arbitrary points is considered in Appendix B.) Besides ellipsoidal coordinates, three other sets of coordinates used for triaxial ellipsoids are: geodetic coordinates $(\phi,\lambda)$ defined by
$\hat{\mathbf{U}}=[\cos\phi\cos\lambda,\cos\phi\sin\lambda,\sin\phi]^{T};$ (A.1a)
parametric coordinates $(\phi^{\prime},\lambda^{\prime})$ defined by
$\mathbf{R}=[a\cos\phi^{\prime}\cos\lambda^{\prime},b\cos\phi^{\prime}\sin\lambda^{\prime},c\sin\phi^{\prime}]^{T};$ (A.1b)
and geocentric coordinates $(\phi^{\prime\prime},\lambda^{\prime\prime})$ defined by
$\hat{\mathbf{R}}=[\cos\phi^{\prime\prime}\cos\lambda^{\prime\prime},\cos\phi^{\prime\prime}\sin\lambda^{\prime\prime},\sin\phi^{\prime\prime}]^{T}.$ (A.1c)
Explicit conversions between any of these coordinates and cartesian coordinates form a common pattern. We obtain cartesian coordinates from geodetic coordinates with
$\mathbf{p}$ $=[a^{n}\cos\phi\cos\lambda,b^{n}\cos\phi\sin\lambda,c^{n}\sin\phi]^{T},$ (A.2a)
$\mathbf{R}$ $=\frac{\mathbf{p}}{\sqrt{p_{x}^{2}/a^{2}+p_{y}^{2}/b^{2}+p_{z}^{2}/c^{2}}},$ (A.2b)
with $n=2$. The opposite conversion is given by
$\mathbf{q}$ $=[X/a^{n},Y/b^{n},Z/c^{n}]^{T},$ (A.3a)
[…]
The corresponding conversions for parametric and geocentric coordinates are given by substituting $n=1$ and $n=0$, respectively, in place of $n=2$. For the parametric conversion, Eq. (A.2b) reduces to $\mathbf{R}=\mathbf{p}$, while for the geocentric conversion, Eq. (A.3a) reduces to $\mathbf{q}=\mathbf{R}$.
Equation (4) defines the conversion from ellipsoidal to cartesian coordinates. This may be inverted with
$\mathbf{q}$ $=[X/a,Y/b,Z/c]^{T},$ (A.4a)
$s$ $=k^{2}q_{x}^{2}+(k^{2}-k^{\prime 2})q_{y}^{2}-k^{\prime 2}q_{z}^{2},$ (A.4b)
$t$ $=\sqrt{s^{2}+4k^{2}k^{\prime 2}q_{y}^{2}},$ (A.4c)
$\cos\beta$ $=\begin{cases}\displaystyle\frac{\sqrt{(t+s)/2}}{k},&\text{if $s\geq 0$},\\ \left|q_{y}/\sin\omega\right|,&\text{otherwise},\end{cases}$ (A.4d)
$\sin\omega$ $=\begin{cases}\displaystyle\operatorname{sign}(q_{y})\frac{\sqrt{(t-s)/2}}{k^{\prime}},&\text{if $s<0$},\\ 0,&\text{if $t=0$},\\ q_{y}/\cos\beta,&\text{otherwise},\end{cases}$ (A.4e)
$\sin\beta$ $=\frac{q_{z}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}},$ (A.4f)
$\cos\omega$ $=\frac{q_{x}}{\sqrt{k^{2}\cos^{2}\beta+k^{\prime 2}}}.$ (A.4g)
For each of the three coordinate systems—geodetic, parametric, and geocentric—a meridian ellipse, defined as a line of constant longitude, lies in a plane containing the $Z$ axis. For geodetic coordinates, this plane is defined by $Y/X=(a/b)^{n}\tan\lambda$, with $n=2$; substitute $n=1$ or $n=0$ for parametric or geocentric coordinates. In general, a “circle of latitude”, a line of constant latitude, only lies in a plane for the parametric latitude (and it is not a circle for $a\neq b$).
Unlike ellipsoidal coordinates, none of these three coordinate systems is orthogonal. However, we can define an azimuth $\zeta$ with respect to “geodetic” north, defined by $\partial\mathbf{R}/\partial\phi$; this north direction is the same for all three systems, namely
$\mathbf{N}^{\prime}=[-XZ/c^{2},-YZ/c^{2},X^{2}/a^{2}+Y^{2}/b^{2}]^{T},$ (A.5)
where the prime is used to distinguish $\mathbf{N}^{\prime}$ from $\mathbf{N}$, which is measured with respect to ellipsoidal coordinates. We can now convert between the cartesian direction $\mathbf{V}$ and either $\alpha$ or $\zeta$.
Appendix B Geodetic coordinates for arbitrary points
Geodetic and ellipsoidal coordinates have natural extensions to arbitrary points in three dimensions. Geodetic coordinates are generalized by giving the height normal to the ellipsoid. Thus, a position is given by
$\mathbf{R}=\mathbf{R}_{0}+h\hat{\mathbf{U}}(\mathbf{R}_{0}),$ (B.1)
where $\mathbf{R}_{0}$ is the closest point on the ellipsoid and $h$ is the height. The full geodetic coordinates are then given by $(\phi,\lambda,h)$.
The extension of ellipsoidal coordinates to three dimensions places an arbitrary point on a confocal ellipsoid defined by
$\frac{X^{2}}{u^{2}+l_{a}^{2}}+\frac{Y^{2}}{u^{2}+l_{b}^{2}}+\frac{Z^{2}}{u^{2}}-1=0,$ (B.2)
where $l_{a}=\sqrt{a^{2}-c^{2}}$ and $l_{b}=\sqrt{b^{2}-c^{2}}$ are linear eccentricities and $u$ is its minor semiaxis. The full ellipsoidal coordinates are $(\beta,\omega,u)$, and the conversion for these to cartesian coordinates is given by Eq. (4), replacing $(a,b,c)$ by $(\sqrt{u^{2}+l_{a}^{2}},\sqrt{u^{2}+l_{b}^{2}},u)$.
We have handled the conversion from $(\phi,\lambda,h)$ and $(\beta,\omega,u)$ to cartesian. Let us address the reverse operation, starting with the conversion of cartesian coordinates to geodetic. This is a standard problem covered, for example, by Bell (1912, §76). The solution is given by finding the largest root $p$ of
$f(p)=\biggl(\frac{aX}{p+l_{a}^{2}}\biggr)^{2}+\biggl(\frac{bY}{p+l_{b}^{2}}\biggr)^{2}+\biggl(\frac{cZ}{p}\biggr)^{2}-1=0.$ (B.3)
Then we have
$\mathbf{R}_{0}$ $=\biggl(\frac{a^{2}X}{p+l_{a}^{2}},\frac{b^{2}Y}{p+l_{b}^{2}},\frac{c^{2}Z}{p}\biggr)^{T},$ (B.4a)
$h$ $=\hat{\mathbf{U}}(\mathbf{R}_{0})\cdot(\mathbf{R}-\mathbf{R}_{0})=(p-c^{2})\,U(\mathbf{R}_{0}),$ (B.4b)
and $(\phi,\lambda)$ are given by Eqs. (A.3).
Ligas (2012) uses Newton’s method to find the root of Eq. (B.3); however, with his choice of starting guess, this sometimes fails to converge. Panou and Korakitis (2022) cure this defect by using the bisection method to find the root. This is guaranteed to converge, but at a high computational cost. Alternatively, Diaz–Toca et al. (2020) use Newton’s method to find the largest root of a sixth-order polynomial obtained by converting Eq. (B.3) to a rational expression; however, Panou and Korakitis (2022) report that this also fails to converge in some instances.
It turns out we can easily fix the problems with Newton’s method applied to Eq. (B.3). First of all, note that $f(p)$ has positive double poles at $p=0$, $l_{b}^{2}$, and $l_{a}^{2}$ and that $f(p)\rightarrow-1$ for $p\rightarrow\pm\infty$. (For now, we assume that $(X,Y,Z)$ are all nonzero.). Therefore, $f(p)=0$ has a unique root for $p\in(0,\infty)$. In this region, $\mathrm{d}f(p)/\mathrm{d}p<0$ and $\mathrm{d}^{2}f(p)/\mathrm{d}p^{2}>0$, and, as a consequence, picking a starting guess for Newton’s method between $p=0$ and the actual root is guaranteed to converge.
To obtain a reasonably tight bound on the root, we note that
$f(p)$ $\leq\frac{c^{2}Z^{2}}{p^{2}}-1,$ (B.5a)
$f(p)$ $\leq\frac{b^{2}Y^{2}+c^{2}Z^{2}}{(p+l_{b}^{2})^{2}}-1,$ (B.5b)
$f(p)$ $\leq\frac{a^{2}X^{2}+b^{2}Y^{2}+c^{2}Z^{2}}{(p+l_{a}^{2})^{2}}-1,$ (B.5c)
$f(p)$ $\geq\frac{a^{2}X^{2}+b^{2}Y^{2}+c^{2}Z^{2}}{p^{2}}-1.$ (B.5d)
Because $\mathrm{d}f(p)/\mathrm{d}p<0$ for $p>0$, this leads to bounds on the positive root, $p_{\mathrm{min}}\leq p\leq p_{\mathrm{max}}$, where
$p_{\mathrm{min}}$ $=\max\Bigl(c\left|Z\right|,\sqrt{b^{2}Y^{2}+c^{2}Z^{2}}-l_{b}^{2},$
$\qquad\qquad\sqrt{a^{2}X^{2}+b^{2}Y^{2}+c^{2}Z^{2}}-l_{a}^{2}\Bigr),$ (B.6a)
$p_{\mathrm{max}}$ $=\sqrt{a^{2}X^{2}+b^{2}Y^{2}+c^{2}Z^{2}}.$ (B.6b)
In implementing Newton’s method, we neglect any term in the definition of $f(p)$ if its numerator vanishes (even though the denominator might also vanish).
Provided that $f(p_{\mathrm{min}})>0$, we can start Newton’s method with $p_{0}=p_{\mathrm{min}}$, and this is guaranteed to converge to the root from below. If $f(p_{\mathrm{min}})\leq 0$, which can only happen if $Z=0$, the required solution is $p=0$. In this case, the expression for $\mathbf{R}_{0}$ is indeterminate (at least one of the components in Eq. (B.4a) involves division by 0), and we proceed as follows:
  • If $X_{0}$ is indeterminate, set $X_{0}=0$ (this can only happen with $X=0$ on a sphere).
  • If $Y_{0}$ is indeterminate, set $Y_{0}=0$ (this can only happen with $Y=0$ on an oblate spheroid).
  • Finally, set $Z_{0}=\pm c\sqrt{1-X^{2}/a^{2}-Y^{2}/b^{2}}$.
This prescription obviates the need to enumerate and treat various subcases as Diaz–Toca et al. (2020); Panou and Korakitis (2022) do.
Turning to the question of converting a cartesian position into ellipsoidal coordinates, we need to find the largest value of $u$ that solves Eq. (B.2). Writing $q=u^{2}$, this becomes the task of finding the largest root of
$g(q)=\frac{X^{2}}{q+l_{a}^{2}}+\frac{Y^{2}}{q+l_{b}^{2}}+\frac{Z^{2}}{q}-1=0.$ (B.7)
The structure of $g(q)$ resembles that of $f(p)$, Eq. (B.3). Since $g(q)$ has 3 simple poles with positive coefficients, there are three real roots, and, because the rightmost pole is at $q=0$ and because $g(q\rightarrow\infty)=-1$, just one of them is positive. As before, bounds can be put on this root $q_{\mathrm{min}}\leq q\leq q_{\mathrm{max}}$, where
$q_{\mathrm{min}}$ $=\max(Z^{2},Y^{2}+Z^{2}-l_{b}^{2},X^{2}+Y^{2}+Z^{2}-l_{a}^{2}),$ (B.8a)
$q_{\mathrm{max}}$ $=X^{2}+Y^{2}+Z^{2}.$ (B.8b)
Provided that $g(q_{\mathrm{min}})>0$, we start Newton’s method with $q_{0}=q_{\mathrm{min}}$, and this converges to the root from below. If $g(q_{\mathrm{min}})\leq 0$ (which can only happen if $Z=0$), the required solution is $q=u=0$.
Of course, we can expand $g(q)$ to obtain a cubic polynomial in $q$, which can be solved analytically. This method is used by Panou and Korakitis (2021). The solution may be subject to unacceptable roundoff error, but it can be refined by using it as the starting point, $q_{0}$, for Newton’s method (which will converge in one or two iterations). In this case, if $g(q_{0})<0$, $q_{1}$ should be replaced by $\max(q_{1},q_{\mathrm{min}})$.
Having determined $u=\sqrt{q}$, $\beta$ and $\omega$ may be found by applying Eqs. (A.4) to the confocal ellipsoid.
Appendix C Newton with bisection in two dimensions
[…]
[scale=0.75]figures/newt2
Figure 7: Bisection in two dimensions. The heavy curves show $f(x,y)=0$ and $g(x,y)=0$. These curves are constrained to lie within the dashed lines, and thus the solution to $f=g=0$ must lie within the parallelogram formed by these lines. The rectangles show bounding boxes for the solution, as explained in the text.
Let us consider the solution of the coupled nonlinear equations
$f(x,y)$ $=f_{x}(x)-f_{y}(y)-f_{0}=0,$ (C.1a)
$g(x,y)$ $=g_{x}(x)+g_{y}(y)-g_{0}=0.$ (C.1b)
These are Eqs. (26) with $x$ standing for $\psi$ or $u$ and $y$ standing for $\theta$ or $v$. Each of $f_{x}$ and $f_{y}$ is an increasing function of its arguments, while $g_{x}$ and $g_{y}$ are non-decreasing functions. There is a unique solution to these equations for given finite $f_{0}$ and $g_{0}$, with the proviso, for umbilical geodesics, that $2b\left|g_{0}\right|$ is less than or equal to the distance between opposite umbilics.
Each of these functions consists of a secular term proportional to $x$ or $y$ and a bounded variation about this term. The slope of the secular part is positive for $f_{x}$ and $f_{y}$ and non-negative for $g_{x}$ and $g_{y}$, and it is straightforward to place an upper bound on the variations (e.g., by summing the magnitudes of the coefficients of the oscillating terms in the Fourier representations). As a consequence, we can bound the solution to a parallelogram in $(x,y)$ space as shown in Fig. 7.
One approach to solving the two-dimensional system is to regard, say, $y$ as the control variable and to write $x=f_{x}^{-1}(f_{y}(y)+f_{0})$. Here, $x$ is found by an inner invocation of Newton’s method. The $g$ equation is solved by an outer invocation. This method is effective but typically requires many function evaluations.
An alternative is to use Newton’s method in two dimensions. This is a straightforward generalization of the normal one-dimensional method, with the reciprocal of the derivative replaced by the inverse of the Jacobian for the system. The parallelogram bound on the solution can be used to estimate a starting point and to detect when the method goes awry. However, in general, it is difficult to establish ever-narrowing bounds on the solution so that a bisection step will provide a good guess, allowing Newton’s method to be continued.
It turns out that for this class of two-dimensional root-finding problems, it is quite easy to establish an axis-aligned bounding rectangle, $x\in[x_{a},x_{b}]$ and $y\in[y_{a},y_{b}]$. This starts by including all of the initial bounding parallelogram—a much looser bound. On each iteration, update one edge of the bounding rectangle depending on the signs of $f$ and $g$. Thus, in the example shown in Fig. 7, the corners of the initial bounding box are marked by circles, and the initial guess, marked by a cross, is at its center. At this point, we find that $f<0$ and $g>0$. From the figure, it is clear that the desired solution lies below this point, and therefore, we update $y_{b}$, giving a new bounding box shown as the light rectangle. This is possible because of the constraints on the slopes of the curves $f=0$ and $g=0$, which, in turn, result from the form of these functions in Eqs. (C.1). With other sign combinations for $f$ and $g$, the other edges of the bounding rectangle can be updated, and in the case where either $f$ or $g$ vanishes, two edges can be updated.
There are several refinements possible. First of all, the method depends on the monotonicity of $f_{x}$, $f_{y}$, $g_{x}$, and $g_{y}$, and, in a numerical context, this is not assured. We remedy this by maintaining a list of all the values of $x\in[x_{a},x_{b}]$ encountered so far and the associated values of $f_{x}(x)$ and $g_{x}(x)$ (and correspondingly for $y$). When a new value of $x$ is inserted, we “clamp” the value of $f_{x}(x)$ and $g_{x}(x)$ to those of the neighboring values.
When a Newton iteration falls outside the bounding box or if the method is converging too slowly, a new starting value is given by the values of $x$ and $y$ that are the midpoints of the largest gaps in their respective lists.
We can be more aggressive in updating the bounding rectangle. Whenever a new $x$ is added to the $x$ list, we update the bounds by checking the signs of $f$ and $g$ with the new $x$ and all the previous $y$ values (and also for each new $y$). This prevents the lists for $x$ and $y$ from growing very long.
Sometimes the procedure given in the previous paragraph leads to a violation of the constraints, which are obvious from Fig. (7): $f(x_{a},y_{b})\leq 0$, $f(x_{b},y_{a})\geq 0$, $g(x_{a},y_{a})\leq 0$, and $g(x_{b},y_{b})\geq 0$. This can happen, for example, if $g=0$ at several of the intersections of the $x$ and $y$ lists. In this case, we update the bounds with only the new value of $(x,y)$.
We also apply this method in the biaxial case for which $g_{y}(y)=0$. This is a degenerate case because it reduces to two one-dimensional root-finding problems: use the $g$ equation Eq. (C.1b) to find $x$ and then, with this value of $x$, use the $f$ equation Eq. (C.1a) to find $y$. The two-dimensional solution described here sometimes suffers from poor convergence because the value of $x$ oscillates between two consecutive floating-point numbers, neither of which exactly satisfies the $g$ equation. In this case, we adjust $g_{0}$ to allow the $g$ equation to be satisfied exactly, and the two-dimensional solution proceeds to solve the $f$ equation in the same manner as the one-dimensional solution.
Appendix D The elliptical billiard problem
[…]
[scale=0.75]figures/similar
Figure 8: A billiard with trajectory marked 11 and 22 (shown as heavy lines) bouncing off the wall of the table at $H$. The foci of the elliptical wall are $F$ and $G$, and standard conventions are used to indicate equal angles. The right triangles $HFF_{1}$ and $HGG_{2}$ are similar, which gives $\left|HF\right|/\left|HG\right|=\left|FF_{1}\right|/\left|GG_{2}\right|$. Also $HFF_{2}$ and $HGG_{1}$ are similar, which gives $\left|HF\right|/\left|HG\right|=\left|FF_{2}\right|/\left|GG_{1}\right|$. The sides of the triangle FGHFGH are $f$, $g$, and $h$.
[…]
Figure 9: Samples of geodesics for an ellipsoid with $a=1.01$, $b=1$, and $c=0$, i.e., an elliptical disk. The geodesics on the upper (resp. lower) face of the ellipse are shown as solid (resp. dashed) lines. Parts (a), (b), and (c) show geodesics with $\gamma>0$, $\gamma=0$, and $\gamma<0$, respectively.
In the limit $c\rightarrow 0$, i.e., $ek\rightarrow 1$, the ellipsoid is flattened to an elliptical disc, and the geodesic problem reduces to the problem of a ball bouncing off the walls of an elliptical billiard table. To make the correspondence exact, we would further (fancifully) stipulate that on each bounce the ball switches between the top and bottom of the table. In this case, the factor $\sqrt{1-e^{2}k^{2}\cos^{2}\beta}$ appearing in Eqs. (17) reduces to $\left|\sin\beta\right|$, and the integrand can be well represented by a Fourier series, provided we split the integral up into pieces depending on the sign of $\sin\beta$; this is the same as the procedure used to obtain the equations for umbilical geodesics Eqs. (27).
But this is a very roundabout way to obtain the trajectory of the billiard ball, which can be obtained by elementary trigonometry. In this case, the simple closed geodesics are: rolling around the edge of the table; following the minor axis of the ellipse (both of these are stable); and bouncing between the foci of the ellipse (this is unstable). These cases correspond, respectively, to geodesics following the major, minor, and median principal ellipses for the ellipsoid.
The conserved quantity $\gamma$ Eq. (14) has a simple interpretation: it is proportional to the product of the angular momenta about the two foci of the ellipse. This conservation law is proved using the similar triangles shown in Fig. 8. Because the speed of the ball is constant, its angular momentum on trajectory $1$ about the focus $F$ is proportional to the perpendicular distance $\left|FF_{1}\right|$, etc. Combining the relationships for the similar triangles in Fig. 8, we obtain
$\left|FF_{1}\right|\left|GG_{1}\right|=\left|FF_{2}\right|\left|GG_{2}\right|,$ (D.1)
which establishes that the product of the angular momenta is conserved on a bounce. Typical paths are shown in Fig. 9; compare these with the geodesics shown in Fig. 2. For $\gamma>0$, Fig. 9(a) (resp. $\gamma<0$, Fig. 9(c)), the path of the billiard ball is tangent to a confocal ellipse (resp. hyperbola).
The instability of the path connecting the two foci, $F$ and $G$, is easily established by considering the triangle $FGH$ in Fig. 8. We adopt the usual nomenclature for triangles, where $F$, $G$, and $H$ measure the interior angles and $f$, $g$, and $h$ are the lengths of the opposite sides. One of Mollweide’s formulas for a triangle gives
$\tan\tfrac{1}{2}F\tan\tfrac{1}{2}G$ $=\frac{(f+g)-h}{(f+g)+h}=\frac{a-\sqrt{a^{2}-b^{2}}}{a+\sqrt{a^{2}-b^{2}}}$
$=\exp(-\Delta),$ (D.2)
where we have substituted $f+g=2a$ and $h=2\sqrt{a^{2}-b^{2}}$, and $\Delta$ is given by either of Eqs. (32) or (39) with $c=0$. On successive bounces, the tangents of the half-angles at the left focus form a geometric progression, increasing by a factor $\exp(2\Delta)$ on each passage through the left focus. If the path is followed in the forward direction, it will coincide with the major axis of the ellipse, while if followed in the reverse direction, it will lie on the major axis in the opposite sense. This is illustrated in Fig. 9(b).
Appendix E Ordinary differential equations for geodesics
The equation for geodesics on a surface is the same as for the motion of a particle constrained to move on the surface but subject to no other forces. The centrifugal acceleration of the particle is $(V^{2}/R_{c})\hat{\mathbf{U}}$, where $R_{c}$ is the radius of curvature in the direction of the velocity $\mathbf{V}$. We will take the speed to be unity (and, of course, the speed is a constant in this problem); thus, time can be replaced by $s$, the displacement along the geodesic, as the independent variable. The differential equations for the geodesic are
$\mathrm{d}\mathbf{R}/\mathrm{d}s$ $=\mathbf{V},$ (E.1a)
$\mathrm{d}\mathbf{V}/\mathrm{d}s$ $=\mathbf{A},$ (E.1b)
where
$\mathbf{A}=-\frac{\mathbf{U}}{U^{2}}\biggl(\frac{V_{x}^{2}}{a^{2}}+\frac{V_{y}^{2}}{b^{2}}+\frac{V_{z}^{2}}{c^{2}}\biggr).$ (E.1c)
This expression for the acceleration $\mathbf{A}$ is obtained by computing the deviation of the particle from the ellipsoid if $\mathbf{V}$ is constant; the acceleration necessary to maintain the particle on the ellipsoid immediately follows. Panou (2013) solves these equations in ellipsoidal coordinates, but this leads to a badly behaved system because of the singular behavior of these coordinates near umbilics. A better approach, adopted by Panou and Korakitis (2019), is to express $\mathbf{R}$ and $\mathbf{V}$ in cartesian coordinates, because there are no singularities in this representation.
Panou and Korakitis (2019) integrated the system using a 4th-order Runge-Kutta scheme. Because this is a relatively low-order method, it’s necessary to use a small step size to control the truncation error. Unfortunately, because a large number of steps are required, this might give an unacceptably large roundoff error. They mitigated this by using “long double” precision (with 64 bits in the fraction as opposed to 53 bits for standard double precision).
A better approach is to use a high-order integration method. Such methods typically adjust the step size automatically to obtain the desired accuracy. After some experimentation, I found satisfactory tools to integrate Eqs. (E.1), as follows:
  • Octave’s ode45: This is an implementation of the Dormand-Prince method.
  • MATLAB’s ode89: A high-order Runge-Kutta method.
  • Boost’s bulirsch_stoer: This Bulirsch-Stoer method uses Richardson extrapolation to obtain an accurate solution, and it can be used with floating-point numbers of any precision.
  • Boost’s bulirsch_stoer_dense_out: This is a variant of the previous method providing “dense output,” i.e., accurately interpolated results within a step.
All these methods allow the direct geodesic problem to be solved about as accurately and about as fast as using Jacobi’s method described in the body of this paper. For example, applying the Bulirsch-Stoer method to the test set for Cayley’s ellipsoid, I find that the average error in the position at point 2 is $120\,\mathrm{ulp}$, and the average CPU time is $15\,\mu\mathrm{s}$. On average, 10 integration steps are required. However, the length of the geodesics in the test set is bounded—they are all “shortest geodesics.” For longer geodesics, the CPU time will scale linearly with distance and the accuracy will degrade.
These errors in the solution given by integrating the ordinary differential equations are an impediment to solving the inverse problem. This depends sensitively on certain properties of the solution, e.g., that a geodesic leaving an umbilic intersects the opposite umbilic. I have stitched up a solution for the inverse problem in the Octave/MATLAB version of GeographicLib by tracking the solution in ellipsoidal coordinates as the solution in cartesian coordinates unfolds (Karney, 2024b). But the code is somewhat messy, because it involves repeated coordinate conversions and has to work around the small, but inevitable, errors in the solution of the direct problem.
Appendix F The stability of closed geodesics
The stability of a geodesic is determined by
$t^{\prime\prime}=-Kt,$ (F.1)
where $t$ is the infinitesimal separation of a geodesic from a reference geodesic, prime indicates differentiation with respect to $s$ (i.e., $t^{\prime\prime}=\mathrm{d}^{2}t/\mathrm{d}s^{2}$), and
$K$ $=\frac{a^{2}c^{2}/b^{6}}{(1+e^{2}k^{\prime 2}\sin^{2}\omega)^{2}(1-e^{2}k^{2}\cos^{2}\beta)^{2}}$
$=\frac{1}{a^{2}b^{2}c^{2}U^{4}}$ (F.2)
is the Gaussian curvature. The first form of $K$ is given by Klingenberg (1982, §3.5.11), and the second is obtained by converting it to cartesian coordinates. Equation (F.1) is solved with two sets of initial conditions,
  • $t(0)=1$, $t^{\prime}(0)=0$, and then $t=M_{12}$ is the forward geodesic scale,
  • $t(0)=0$, $t^{\prime}(0)=1$, and then $t=m_{12}$ is the reduced length, and $t^{\prime}=M_{21}$ is the reverse geodesic scale.
Some of the properties of $m_{12}$, $M_{12}$, and $M_{21}$ are given in Karney (2013, §3). I solve for these quantities by supplementing the ordinary differential equations in Appendix E with Eq. (F.1). This gives a system of ten first-order differential equations, six for the geodesic and four for $m_{12}$ and $M_{12}$.
For a closed geodesic on one of the principal ellipses, $K(s)$ is a periodic function with period $s_{0}$, one half of the perimeter of the ellipse. Thus, Eq. (F.1) is an example of Hill’s equation (Olver et al., 2010, §28.29), which can be solved using Floquet’s theorem. To determine the behavior of the equation for large $s$, it suffices to solve it over one period and form the matrix,
$\mathcal{M}=\begin{bmatrix}M_{12}&m_{12}\\ M^{\prime}_{12}&m^{\prime}_{12}\end{bmatrix}=\begin{bmatrix}M_{12}&m_{12}\\ {\displaystyle-\frac{1-M_{12}M_{21}}{m_{12}}}&M_{21}\end{bmatrix}$, (F.3)
where all the terms in the matrix are evaluated at $s=s_{0}$. The Wronskian of this system is unity, $\det\mathcal{M}=1$. If the initial conditions are $t(0)=W_{0}$ and $t^{\prime}(0)=w_{0}$, then, after an integer $l$ periods, we have $t(ls_{0})=W_{l}$ and $t^{\prime}(ls_{0})=w_{l}$ where
$\begin{bmatrix}W_{l}\\ w_{l}\end{bmatrix}=\mathcal{M}^{l}\begin{bmatrix}W_{0}\\ w_{0}\end{bmatrix}$. (F.4)
The stability of the system is determined by the eigenvalues of $\mathcal{M}$. Writing $\operatorname{tr}\mathcal{M}=M_{12}+M_{21}=2M$, the eigenvalues are
$\lambda_{1,2}=M\pm\sqrt{M^{2}-1},$ (F.5)
with $\lambda_{1}\lambda_{2}=1$. We distinguish four cases:
  • if $\left|M\right|>1$, one of $\left|\lambda_{1,2}\right|$ is greater than unity, and the solution is exponentially unstable;
  • if $\left|M\right|<1$, $\left|\lambda_{1,2}\right|$ are both unity, and the solution is stable (bounded oscillations);
  • if $\left|M\right|=1$ and the off-diagonal terms of $\mathcal{M}$ vanish, the solution is stable;
  • if $\left|M\right|=1$ and at least one off-diagonal term of $\mathcal{M}$ is nonzero, the solution is linearly unstable.
Computing $M$ for the principal ellipses of Cayley’s ellipsoid, we have M=0.3634M=-0.3634, 1.6399-1.6399, and $0.4274$ for the minor, median, and major ellipses. This confirms that the median ellipse is exponentially unstable, while the other two are stable. Examples where $\left|M\right|=1$ are: great circles on a sphere (stable), the equator for an oblate ellipsoid with integer $b/c$ (stable), and meridian ellipses on biaxial ellipsoids (linearly unstable).
The rate of instability for the median ellipse is reflected in the quantity $\Delta$ introduced in Sec. 6. However, we can generalize that result to apply to any of the principal ellipses. In Hart’s equation for $\Delta$ Eq. (39), $a$ and $c$, the semiaxes of the median ellipse, appear symmetrically, while $b$ occupies a privileged position. By making suitable exchanges between $a$, $b$, and $c$, we obtain the corresponding values of $\Delta$ for all three principal ellipses, giving $\Delta=1.1989\,i$, $1.0783$, and 2.0125i2.0125\,i (with $i=\sqrt{-1}$). We now find that $M$ and $\Delta$ are related by $M=-\cosh\Delta$.
Data availability
Test data is available at Karney (2024c). Additional test data is available upon request.
Code availability
A C++ implementation of the algorithms given in this paper is given in GeographicLib, version 2.6, available at https://github.com/geographiclib/geographiclib/releases/tag/r2.6. Version 2.7 of GeographicLib will be released soon; this will fix bugs found during the preparation of this manuscript.

### v2

@@
 1 Introduction
-The geodesic is the shortest path between two points on a surface, and it plays a crucial role in geodesy, where the earth is typically modeled as an ellipsoid of revolution, a biaxial ellipsoid.
+The shortest path between two points on a surface is a geodesic, and it plays a crucial role in geodesy, where the earth is typically modeled as an ellipsoid of revolution, a biaxial ellipsoid.
 The main geodesic problems are (1) given a starting point and a direction, find the point a certain distance away, the “direct” geodesic problem, and (2) to find the length and direction of the geodesic connecting two points, the “inverse” geodesic problem.
-The path of a geodesic is also given by the motion of a point mass sliding on the surface of the ellipsoid without friction and in the absence of external forces.
-This allows geodesics to be extended indefinitely.
-Recently, there has been interest in using a refined model of the earth as a triaxial ellipsoid (Panou et al., 2020).
-Consequently, there is interest in obtaining reliable solutions to the geodesic problem in this case.
+The path of a geodesic is also given by a point mass constrained to the surface of the ellipsoid and moving without friction and in the absence of external forces—this path is “as straight as possible.” This allows geodesics to be extended indefinitely.
+Because of their central role in geodesy, there is naturally an interest in exploring the behavior of geodesics when the earth is modeled as a triaxial ellipsoid (Burša and Fialová, 1993).
+In addition, there are other bodies in our solar system which are not rotationally symmetric but which can be approximated by a triaxial ellipsoid.
+Consequently, reliable solutions to the geodesic problems for a triaxial ellipsoid would be useful.
 The solution of the direct geodesic problem in the biaxial case is relatively straightforward.
@@
 It therefore came as a surprise when Jacobi (1839) found that the geodesic problem could be reduced to quadrature in this case too, with the solution given in terms of one-dimensional integrals.
-(We know the date, given in the abstract, for this discovery, because of a letter he wrote on December 28, 1838, to F.
+(We know the date, given in the abstract, for this discovery, because of a letter he wrote two days later to F.
 W.
 Bessel, his neighbor in Königsberg.) Jacobi (1843, §28) expanded on his method in his Lectures on Dynamics, and the result was generalized by Liouville (1846, §§20–21) to apply to so-called Liouville surfaces.
-The qualitative properties of the solution can be found in several textbooks (Darboux, 1894; Hilbert and Cohn-Vossen, 1952; Klingenberg, 1982; Arnol’d, 1989; Berger, 2010).
+The qualitative properties of the solution can be found in several textbooks (Darboux, 1894; Hilbert and Cohn-Vossen, 1952; Klingenberg, 1982; Arnold, 1989; Berger, 2010).
 However, unlike the case of the biaxial ellipsoid, where Bessel (1825) provided a prescription for computing geodesics, little effort was given to implementing Jacobi’s solution.
@@
 2 Ellipsoidal coordinates
-[…]
-[scale=0.75]figures/graticule
-Figure 1: The ellipsoidal grid showing lines of constant $\beta$ and $\omega$.
-The grid spacing is $15^{\circ}$.
-The heavy lines show the minor ($X=0$ or $\cos\omega=0$), median ($Y=0$ or $\cos\beta\sin\omega=0$), and major ($Z=0$ or $\sin\beta=0$) principal ellipses of the ellipsoid.
-The parameters of the ellipsoid are $a=1.01$, $b=1$, and $c=0.8$, and it is viewed in an orthographic projection looking at the point with geodetic coordinates $\phi=40^{\circ}$, $\lambda=30^{\circ}$.
 Jacobi’s insight was to express the equations of the geodesic in terms of ellipsoidal coordinates; this allows the equations to be reduced to one-dimensional integrals through the separation of variables.
@@
 $\mathbf{R}=\begin{bmatrix}a\cos\omega\sqrt{k^{2}\cos^{2}\beta+k^{\prime 2}}\\ b\cos\beta\sin\omega\\ c\sin\beta\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\end{bmatrix}$. (4)
+Figure 1: The ellipsoidal grid showing lines of constant $\beta$ and $\omega$.
+The grid spacing is $15^{\circ}$.
+The heavy lines show the minor ($X=0$ or $\cos\omega=0$), median ($Y=0$ or $\cos\beta\sin\omega=0$), and major ($Z=0$ or $\sin\beta=0$) principal ellipses of the ellipsoid.
+The parameters of the ellipsoid are $a=1.01$, $b=1$, and $c=0.8$, and it is viewed in an orthographic projection looking at the point with geodetic coordinates $\phi=40^{\circ}$, $\lambda=30^{\circ}$.
 Lines of constant $\beta$ and $\omega$ define a grid, illustrated in Fig.
@@
 Furthermore, the direction of a geodesic is
-$\mathbf{V}=\sin\alpha\hat{\mathbf{E}}+\cos\alpha\hat{\mathbf{N}}$, (7)
-where $\alpha$ is the azimuth of the geodesic measured clockwise from a line of constant $\omega$. At the pole of an oblate ellipsoid, we take the limit $\cos\beta\rightarrow 0+$, to give
+$\mathbf{V}=\hat{\mathbf{E}}\sin\alpha+\hat{\mathbf{N}}\cos\alpha,$ (7)
+where $\alpha$ is the azimuth of the geodesic measured clockwise from a line of constant $\omega$ and the hat symbol denotes a unit vector.
+At the pole of an oblate ellipsoid, we take the limit $\cos\beta\rightarrow 0+$, to give
 $\hat{\mathbf{E}}$ $=[-\sin\omega,\cos\omega,0]^{T},$ (8a)
@@
 this follows from setting $\gamma=0$ in Eq. (14), given below. Expanding $\mathbf{R}$ about an umbilic to second order in $\cos\beta$ and $\sin\omega$, we find
-[…]
+$\mathbf{V}=-\sin\beta(\hat{\mathbf{E}}\sin(2\alpha)-\hat{\mathbf{N}}\cos(2\alpha)),$ (12)
 where we have chosen the sign in Eq. (11) as $\pm 1=-\sin\beta\cos\omega$ to yield the normal convention that $\alpha$ measures angles clockwise.
@@
 3 Qualitative behavior
-[…]
-[…]
-[…]
-[scale=0.75]figures/geodesicc
-[…]
-[…]
 Figure 2: Samples of geodesics on an ellipsoid with the same parameters and viewpoint as Fig.
@@
 For a biaxial ellipsoid, the equator and all the meridians are simple (not self-intersecting) closed geodesics.
-On the other hand, for a triaxial ellipsoid (provided it is not too eccentric), there are only three simple closed geodesics, namely the 3 principal ellipses.
-The major and minor ellipses are stable; if they are perturbed, the resulting geodesic oscillates about the original ellipse.
+On the other hand, for a triaxial ellipsoid, provided it is not too eccentric, there are only three simple closed geodesics, namely the 3 principal ellipses.
+(For sufficiently eccentric ellipsoids, $e^{2}k^{3/2}>3/4$, other simple closed geodesics that oscillate about the major principal ellipse are possible.) The major and minor ellipses are stable; if they are perturbed, the resulting geodesic oscillates about the original ellipse.
 However, the median ellipse is unstable; if the geodesic is perturbed, it swings away from the $Y=0$ plane before returning to the original ellipse, but now traveling along it in the opposite direction.
@@
 4 Jacobi’s solution
-[…]
-[scale=0.75]figures/cayley
-Figure 3: The graphical method for plotting umbilical geodesics given by Cayley (1872).
-The semiaxes of the ellipsoid are $a=\sqrt{1000}$, $b=\sqrt{500}$, $c=\sqrt{250}$, and it is viewed here in an orthogonal projection along the $Y$ axis.
-The dotted lines are lines of $\beta=\beta^{(f)}_{n}$ (going bottom to top) or $\omega=\omega^{(f)}_{n}$ (going left to right), as defined in the text, with the separation constant given by $\Delta^{(f)}=1/\sqrt{160}$ (matching Cayley’s choice).
-In this figure, the lines of constant $\beta$ and $\omega$ are labeled with the corresponding values of $n$ (these values are given in Table 2).
-The geodesics, shown as heavy lines, connect the vertices of the resulting mesh, and all converge on the umbilic labeled $U$, or on the neighboring one (not shown in the figure).
-[…]
-[scale=0.75]figures/cayleydist
-Figure 4: Marking the distance along geodesics.
-The heavy lines are those geodesics in Fig.
-3 that converge on the umbilic $U$.
-As in that figure, the dotted lines are lines of constant $\beta$ or $\omega$; however, in this case, the values are given by equal increments of $\Delta^{(g)}=1/10$ in the distance functions (these values are given in Table 2).
-The dashed lines connect the vertices of the mesh, and these mark off distance intervals of $b/10$ along the geodesics.
 Here we summarize the solution of the geodesic problem following Darboux (1894, §§583–585); a comparable treatment is given by Klingenberg (1982, §§3.5.4–3.5.6).
@@
 and the distance $s$ along the geodesic is given by Darboux’s Eq. ($33^{\prime}$),
-$\frac{s+s_{1}}{b}$ $=\int\frac{k\cos^{2}\beta\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
-$\quad{}\pm\int\frac{k^{\prime}\sin^{2}\omega\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega.$ (17b)
+$\frac{s+s_{1}}{b}$ $=\int\frac{k^{2}\cos^{2}\beta\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
+$\quad{}\pm\int\frac{k^{\prime 2}\sin^{2}\omega\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega.$ (17b)
 Here $\delta$ and $s_{1}$ are constants given by the initial conditions.
@@
 In the following, we drop the parametric arguments for the $f$ and $g$ functions; $\kappa,\epsilon,\mu]=[k^{2},e^{2},-\gamma]$ are implied for functions of $\beta$, and $k^{\prime 2},-e^{2},\gamma]$ for functions of $\omega$.
+Figure 3: The graphical method for plotting umbilical geodesics given by Cayley (1872).
+The semiaxes of the ellipsoid are $a=\sqrt{1000}$, $b=\sqrt{500}$, $c=\sqrt{250}$, and it is viewed here in an orthogonal projection along the $Y$ axis.
+The dotted lines are lines of $\beta=\beta^{(f)}_{n}$ (going bottom to top) or $\omega=\omega^{(f)}_{n}$ (going left to right), as defined in the text, with the separation constant given by $\Delta^{(f)}=1/\sqrt{160}$ (matching Cayley’s choice).
+In this figure, the lines of constant $\beta$ and $\omega$ are labeled with the corresponding values of $n$ (these values are given in Table 2).
+The geodesics, shown as heavy lines, connect the vertices of the resulting mesh, and all converge on the umbilic labeled $U$, or on the neighboring one (not shown in the figure).
+Figure 4: Marking the distance along geodesics.
+The heavy lines are those geodesics in Fig.
+3 that converge on the umbilic $U$.
+As in that figure, the dotted lines are lines of constant $\beta$ or $\omega$; however, in this case, the values are given by equal increments of $\Delta^{(g)}=1/10$ in the distance functions (these values are given in Table 2).
+The dashed lines connect the vertices of the mesh, and these mark off distance intervals of $b/10$ along the geodesics.
 Table 2: Values of $\beta^{(f)}_{n}$, $\omega^{(f)}_{n}$, $\beta^{(g)}_{n}$, and $\omega^{(g)}_{n}$ used for the coordinate meshes in Figs.
@@
 (17) is singular and that, if $\gamma=0$, the integrals themselves are singular.
-In this section, we address the circum- and transpolar cases where $\gamma\neq 0$.
+In this section, we address the circumpolar and transpolar cases where $\gamma\neq 0$.
 We deal with these cases together by denoting the rotating and librating coordinates as $\theta$ and $\phi$, respectively.
@@
 Geodesics on a biaxial ellipsoid are well understood. However, it’s instructive to see how we can recover the biaxial solution from Jacobi’s.
-For biaxial ellipsoids, the functions $f_{\theta}$ and $g_{\theta}$ are evaluated with $\kappa=0$ to give
+For biaxial ellipsoids, we substitute $\kappa=0$ in Eqs. (21) to give
 $f_{\theta}(\theta;\kappa=0)$ $=\frac{\theta}{\sqrt{\left|\gamma\right|}},$ (42a)
 $g_{\theta}(\theta;\kappa=0)$ $=0.$ (42b)
-The functions $f_{\psi}$ and $g_{\psi}$, with $\kappa=1$, become
+Similarly, we substitute $\kappa=1$ in Eqs. (25) to obtain
 $f_{\psi}(\psi;\kappa=1)$ $=\int_{0}\frac{\sqrt{1-\epsilon(\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi)}}{\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi}\,\mathrm{d}\psi,$ (43a)
 $g_{\psi}(\psi;\kappa=1)$ $=\int_{0}\sqrt{1-\epsilon(\cos^{2}\psi+\left|\gamma\right|\sin^{2}\psi)}\,\mathrm{d}\psi.$ (43b)
+In both Eqs.
+(42) and (43), I have replaced $\left|\mu\right|$ by $\left|\gamma\right|$ in order to facilitate their substitution into Eqs.
+(26).
 In the spherical limit, we carry out the integration to give
@@
 10 The inverse problem
-[…]
-[scale=0.75]figures/cuta
-[…]
-[scale=0.75]figures/cutb
 Figure 5: Geodesics emanating from a single point on an ellipsoid.
@@
 Part (b) shows the geodesics from (a) as dotted lines, and these are continued (as solid lines) until they meet at the line of curvature $\omega=\omega_{1}+180^{\circ}$ (shown as a heavy line).
-[…]
-[scale=0.75]figures/cutunroll
-[…]
-[scale=0.75]figures/hybrid
 Figure 6: (a) The geodesics in Fig.
@@
 I know of no corresponding formula for $m_{12}$ for the triaxial case.
-Nevertheless, it is possible to find the first conjugate point by the condition that $\beta$ has completed half an oscillation, which in turn implies that $\psi$ has advanced by $\pi$.
+Nevertheless, for geodesics with $\alpha_{1}=\pm{\textstyle\frac{1}{2}}\pi$, the first conjugate point is given by the condition that $\beta$ has completed half an oscillation, which in turn implies that $\psi$ has advanced by $\pi$.
 Equation (26a) can be used to give the value of $\omega$ at that point.
@@
 This is known as Ivory’s Lemma.
-11 Implementation
-An implementation of the solutions of the direct and inverse geodesic problems is provided with version 2.6 of the C++ library, GeographicLib (Karney, 2025).
+11 Implementation and results
+An implementation of the solutions of the direct and inverse geodesic problems is provided with version 2.7 of the C++ library, GeographicLib (Karney, 2025a).
 It’s practically impossible to exhaustively document the methods here; the reader is referred to the code for details.
@@
 10.
-The test set (Karney, 2024c) contains 500 000500\,000 geodesics, computed at high precision, with the coordinates of the endpoints given as integer degrees.
+The test set (Karney, 2024b) contains 500 000500\,000 geodesics, computed at high precision, with the coordinates of the endpoints given as integer degrees.
 My testing also included other ellipsoids with $a/c=2$, including oblate and prolate ellipsoids.
-In addition, I tested with spheres ($a=c$) with different values of $k$, and with a triaxial model of the earth (Panou et al., 2020).
+In addition, I tested with spheres ($a=c$) with different values of $k$, and with a triaxial model of the earth (Burša and Fialová, 1993).
 In the following, I will only report results for the published dataset for Cayley’s ellipsoid.
@@
 An implementation of the coordinate conversions given in Appendix A is included in GeographicLib, as is sample code for solving the ordinary differential equations for the direct geodesic problem (Appendix E).
-12 Discussion
+12 Conclusions
 In this paper, I have described the implementation of Jacobi’s solution to the direct geodesic problem for a triaxial ellipsoid.
@@
 Appendix C Newton with bisection in two dimensions
-[…]
-[scale=0.75]figures/newt2
 Figure 7: Bisection in two dimensions.
@@
 However, in general, it is difficult to establish ever-narrowing bounds on the solution so that a bisection step will provide a good guess, allowing Newton’s method to be continued.
-It turns out that for this class of two-dimensional root-finding problems, it is quite easy to establish an axis-aligned bounding rectangle, $x\in[x_{a},x_{b}]$ and $y\in[y_{a},y_{b}]$.
+It turns out that for the class of two-dimensional root-finding problems given by Eq.
+(C.1), it is quite easy to establish an axis-aligned bounding rectangle, $x\in[x_{a},x_{b}]$ and $y\in[y_{a},y_{b}]$.
 This starts by including all of the initial bounding parallelogram—a much looser bound.
@@
 Appendix D The elliptical billiard problem
-[…]
-[scale=0.75]figures/similar
 Figure 8: A billiard with trajectory marked 11 and 22 (shown as heavy lines) bouncing off the wall of the table at $H$.
@@
 The sides of the triangle FGHFGH are $f$, $g$, and $h$.
-[…]
-[…]
 Figure 9: Samples of geodesics for an ellipsoid with $a=1.01$, $b=1$, and $c=0$, i.e., an elliptical disk.
@@
 Appendix E Ordinary differential equations for geodesics
-The equation for geodesics on a surface is the same as for the motion of a particle constrained to move on the surface but subject to no other forces.
+Laplace (1829, §8) showed that a particle constrained to move on a surface but subject to no other forces follows a geodesic.
 The centrifugal acceleration of the particle is $(V^{2}/R_{c})\hat{\mathbf{U}}$, where $R_{c}$ is the radius of curvature in the direction of the velocity $\mathbf{V}$.
-We will take the speed to be unity (and, of course, the speed is a constant in this problem); thus, time can be replaced by $s$, the displacement along the geodesic, as the independent variable.
+We will take the speed $V$ to be unity (and, of course, the speed is a constant in this problem); thus, time can be replaced by $s$, the displacement along the geodesic, as the independent variable.
 The differential equations for the geodesic are
@@
 This depends sensitively on certain properties of the solution, e.g., that a geodesic leaving an umbilic intersects the opposite umbilic.
-I have stitched up a solution for the inverse problem in the Octave/MATLAB version of GeographicLib by tracking the solution in ellipsoidal coordinates as the solution in cartesian coordinates unfolds (Karney, 2024b).
+I have stitched up a solution for the inverse problem in the Octave/MATLAB version of GeographicLib by tracking the solution in ellipsoidal coordinates as the solution in cartesian coordinates unfolds (Karney, 2025b).
 But the code is somewhat messy, because it involves repeated coordinate conversions and has to work around the small, but inevitable, errors in the solution of the direct problem.
@@
 Data availability
-Test data is available at Karney (2024c). Additional test data is available upon request.
+Test data is available at Karney (2024b). Additional test data is available upon request.
 Code availability
-A C++ implementation of the algorithms given in this paper is given in GeographicLib, version 2.6, available at https://github.com/geographiclib/geographiclib/releases/tag/r2.6.
-Version 2.7 of GeographicLib will be released soon; this will fix bugs found during the preparation of this manuscript.
+The C++ implementation of the algorithms given in this paper is given in GeographicLib, version 2.7, available at https://github.com/geographiclib/geographiclib/releases/tag/r2.7.
+This version is archived at https://doi.org/10.5281/zenodo.18011538.